### v2

# 2cm2.5cm2.5cm2cm

Abstract
We propose a general class of INteger-valued Generalized AutoRegressive Conditionally Heteroscedastic (INGARCH) processes by allowing time-varying mean and dispersion parameters, which we call time-varying dispersion INGARCH (tv-DINGARCH) models. More specifically, we consider mixed Poisson INGARCH models and allow for dynamic modeling of the dispersion parameter (as well as the mean), similar to the spirit of the ordinary GARCH models. We derive conditions to obtain first and second-order stationarity, and ergodicity as well. Estimation of the parameters is addressed and their associated asymptotic properties are established as well. A restricted bootstrap procedure is proposed for testing constant dispersion against time-varying dispersion. Monte Carlo simulation studies are presented for checking point estimation, standard errors, and the performance of the restricted bootstrap approach. We apply the tv-DINGARCH process to model the weekly number of reported measles infections in North Rhine-Westphalia, Germany, from January 2001 to May 2013, and compare its performance to the ordinary INGARCH approach.
Keywords: Autocorrelation; Count time series; Overdispersion; Time-varying dispersion parameter; Volatility.
1 Introduction
Modeling count time series data is a challenging and very exciting research topic with applications in many different areas such as epidemiology, sociology, economics, and health science. A well-established methodology, for dealing with count time series data, has been developed under the framework of INteger-valued Generalized AutoRegressive Conditional Heteroscedastic (INGARCH) models, which have been initially studied and explored by Heinen (2003), Ferland et al. (2006), Fokianos et al. (2009), and Fokianos (2011). The INGARCH nomenclature emerges from the fact that for a Poisson distribution (this assumption is imposed in the aforementioned papers), the mean equals its variance, therefore modeling of mean implies modeling of variance like the classic continuous GARCH models introduced by Bollerslev (1986). Consequently, such models are considered in the literature as integer counterpart of the GARCH models although this terminology should be used cautiously because ordinary GARCH processes do not consider dynamics for the conditional mean; for instance, see Fokianos et al. (2009).
A Poisson $\mbox{INGARCH}(p,q)$ (with $p,q\in\mathbb{N}_{0}$) model $\{Y_{t}\}_{t\in\mathbb{N}}$ is defined by
$Y_{t}|\mathcal{F}_{t-1}\sim\mbox{Poisson}(\lambda_{t}),\quad\lambda_{t}\equiv E(Y_{t}|\mathcal{F}_{t-1})=\beta_{0}+\sum_{i=1}^{p}\beta_{i}Y_{t-i}+\sum_{j=1}^{q}\alpha_{j}\lambda_{t-j},$ (1)
where $\mathcal{F}_{t-1}=\sigma\{Y_{t-1},\ldots,Y_{0},\lambda_{0}\}$, $\beta_{0}>0$, $\beta_{i}\geq 0$ and $\alpha_{j}\geq 0$, for all $i=1,\ldots,p$ and $j=1,\ldots,q$. Since $E(Y_{t}|\mathcal{F}_{t-1})=\mbox{Var}(Y_{t}|\mathcal{F}_{t-1})=\lambda_{t}$, the model specifies dynamic processes for both the conditional mean and variance, as previously stated.
The Poisson INGARCH model has been extensively considered in the literature. However, due to limitations for fitting adequately real count time series data (e.g. not capturing all the sources of overdispersion although the model itself is overdispersed, inflation or deflation of zeros), other variants have been proposed based on such as the negative binomial INGARCH process by Zhu (2011) and Christou and Fokianos (2014) or the more general mixed Poisson INGARCH models by Christou and Fokianos (2015), Silva and Barreto-Souza (2019), among others. Zero-inflated versions of the Poisson and negative binomial INGARCH models were proposed by Zhu (2012a), while processes dealing with both overdispersion and underdispersion were proposed by Zhu (2012b, c) and Xu et al. (2012).
An alternative process to the linear model for the conditional mean, as in Equation (1), was introduced by Fokianos and Tjøstheim (2011) through the log-linear INGARCH processes. These models cope with both negative and positive autocorrelation function and allow inclusion of covariates in a straightforward way. Another different approach was recently proposed by Weiß et al. (2022), where a softplus link function is assumed instead of a logarithmic one.
Based on the discussion so far, the Poisson and Negative Binomial distributions are frequently used in applications. In several cases, the Negative Binomial model fits more adequate due to the fact that this distribution depends on a dispersion parameter which facilitates flexible modeling. The dispersion parameter is usually assumed fixed and is estimated by using the Pearson residuals; see Christou and Fokianos (2014) for more. The main goal of this paper is to propose a novel and general class of INGARCH processes based on the mixed Poisson distributions (the Negative Binomial distribution falls in this class) by allowing time-varying mean and dispersion parameters. We call these models time-varying dispersion INGARCH (tv-DINGARCH) models. Hence, we develop methodology and study new models for integer-valued time series, where the assumption of constant dispersion might not be hold. The advantage of the tv-DINGARCH processes over the ordinary INGARCH models will be illustrated in Section 5, where we show that such processes fit better real data whose stylistic facts cannot be dealt with the existing INGARCH models.
For example, in a count regression context, Barreto-Souza and Simas (2016) demonstrated through a data application on the attendance behavior of high school juniors that a constant dispersion assumption can be violated. Generalized linear models allowing regression structures for both mean and dispersion/precision parameters have been discussed previously by Efron (1986) and Smyth (1989), among others. In this work, we develop further this line of research by studying in detail joint modeling of mean and dispersion parameters in the context of dependent data. We put special emphasis in the case of INGARCH models for count time series and demonstrate the usefulness of such methodological development. Besides exploring the traditional INGARCH models, the proposed class contains the case that admits constant mean and time-dependent variance, similar to the spirit of the ordinary GARCH models. Another important feature is volatility modeling, which is a well-explored topic for the case of continuous valued time series but has been neglected in the context of integer-valued time series. Although the traditional Poisson INGARCH processes consider a time-dependent conditional variance, this in turn is driven by the dynamics of the mean process. The main disadvantage of this approach is that imposes severe restrictions as illustrated in the real data application in this paper. The tv-DINGARCH models relax such an assumption by considering a time-dependent dispersion parameter, therefore also controlling the conditional variance and allowing for additional source of volatility.
In the context of time-varying models for INGARCH processes, Roy and Karmakar (2021) introduced a Poisson INGARCH model with time-varying coefficients and Ratnayake and Samaranayake (2023) study a time-varying INGARCH model based on the zero-inflated Poisson distribution. These authors consider dynamics for the inflation parameter which is assumed to depend on exogenous variables. Furthermore, Doukhan et al. (2022) study in detail first order non-stationary INGARCH models and prove mixing conditions under natural assumptions. Our approach is different in several aspects. First, we consider mixed Poisson distributions (rather than Poisson and zero-inflated Poisson models) which cover a broad class of count distributions that include the Poisson and Negative Binomial. In addition, we consider dynamics that enter through the dispersion parameter. One advantage of the latter feature is that we are able to establish desirable stability properties for the models we consider, such as stationarity and ergodicity, based on e-chain theory.
The paper is organized as follows. In Section 2, we define the class of tv-DINGARCH models and then derive its stochastic properties. We establish conditions ensuring stationarity and ergodicity of the count processes. Section 3 is devoted to the estimation of the parameters and the derivation of their associated asymptotic properties as well. Aiming at testing constant dispersion in practice, a restricted bootstrap procedure is proposed in Section 4. Monte Carlo simulation studies are presented for checking point estimation, standard errors, and the performance of the restricted bootstrap approach. In Section 5, we apply the tv-DINGARCH process to model the weekly number of reported measles infections in North Rhine-Westphalia, Germany, from January 2001 to May 2013, and compare its performance to the ordinary INGARCH approach. Some technical results and proofs are provided in the Appendix. This paper contains Supplementary Material, which is available upon request.
Notation:
We say that a random variable $Y$ follows a mixed Poisson (MP) distribution if satisfies the stochastic representation $Y|Z=z\sim\mbox{Poisson}(\lambda z)$, with $Z$ following some non-negative distribution with $E(Z)=1$ (standardization) and $\mbox{Var}(Z)=\phi^{-1}$, for $\lambda,\phi>0$. We denote $Y\sim\mbox{MP}(\lambda,\phi)$. In this case, $E(Y)=\lambda$ and $\mbox{Var}(Y)=\lambda+\lambda^{2}/\phi$. A random variable $Z$ following a Gamma distribution, with respective shape and scale parameters $a>0$ and $b>0$, is denoted by $Z\sim\mbox{Gamma}(a,b)$, where $E(Z)=a/b$ and $\mbox{Var}(Z)=a/b^{2}$. If the mixing variable $Z\sim\mbox{Gamma}(\phi,\phi)$, then we obtain that $Y$ follows a negative binomial (NB) distribution, i.e. $Y\sim\mbox{NB}(\lambda,\phi)$.
For an $l$-dimensional vector ${\bf x}=(x_{1},\ldots,x_{l})^{\top}$, let $\|{\bf x}\|_{p}=(\sum_{i=1}^{l}|x_{i}|^{p})^{1/p}$, for $p\in[1,\infty)$, and $\|{\bf x}\|_{p}=\max_{1\leq i\leq l}|x_{i}|$ for $p=\infty$. The induced $p$-norm for a $m\times l$ matrix ${\bf C}$ is then defined by $\|{\bf C}\|_{p}=\max_{{\bf x}\neq{\bf 0}}\{\|{\bf Cx}\|_{p}/\|{\bf x}\|_{p}:\,{\bf x}\in\mathbb{R}^{l}\}$, for $p\in[1,\infty]$.
2 The tv-DINGARCH Models
In this section we propose a class of time-varying dispersion INGARCH models by allowing both the mean and the dispersion parameter to depend on lagged valued of the observed series and their past values as follows.
Definition 2.1.
[…]
$\lambda_{t}=f(Y_{t-1},\ldots,Y_{t-p_{1}};\lambda_{t-1},\ldots,\lambda_{t-q_{1}}),\quad\phi_{t}=g(Y_{t-1},\ldots,Y_{t-p_{2}};\phi_{t-1},\ldots,\phi_{t-q_{2}}),$ (2)
where $\mathcal{F}_{t-1}=\sigma\{Y_{t-1},\ldots,Y_{0},\phi_{0},\lambda_{0}\}$, and $(\lambda_{0},\phi_{0})$ denoting some starting value, $f:\mathbb{N}^{p_{1}}\times(0,\infty)^{q_{1}}\rightarrow(0,\infty)$, and $g:\mathbb{N}^{p_{2}}\times(0,\infty)^{q_{2}}\rightarrow(0,\infty)$.
Of particular interest to our study and to diverse applications is the negative binomial (NB) model when it is imposed in Definition 2.1. Then, the conditional probability function of $Y_{t}$ given $\mathcal{F}_{t-1}$ is given by
$P(Y_{t}=y|\mathcal{F}_{t-1})=\dfrac{\Gamma(y+\phi_{t})}{y!\Gamma(\phi_{t})}\left(\dfrac{\lambda_{t}}{\lambda_{t}+\phi_{t}}\right)^{y}\left(\dfrac{\phi_{t}}{\lambda_{t}+\phi_{t}}\right)^{\phi_{t}},\quad y\in\mathbb{N}_{0}.$
The negative binomial model will be assumed to prove finiteness of moments and to derive the asymptotic theory for conditional maximum likelihood estimators in what follows. Some of results hold in general for mixed Poisson models, see Theorem 2.1, which states stationarity and ergodicity of time-varying dispersion INGARCH models.
Definition 2.1 is general and some additional assumptions on the function $f$ and $g$ are necessary to study in detail these processes. We consider the first order model ($p_{1}=p_{2}=q_{1}=q_{2}=1$) linear tv-DINGARCH processes, which will be defined in what follows. This particular linear parametric form is a common choice considered in the literature but also enables a thorough study of stability properties of processes defined by Definition 2.1.
Definition 2.2.
[…]
$\lambda_{t}=\beta_{0}+\beta_{1}Y_{t-1}+\beta_{2}\lambda_{t-1},\quad\phi_{t}=\alpha_{0}+\alpha_{1}Y_{t-1}+\alpha_{2}\phi_{t-1},$ (3)
where $\beta_{0},\alpha_{0}>0$ and $\beta_{1},\beta_{2},\alpha_{1},\alpha_{2}\geq 0$.
Remark 2.1.
Note that we are using $Y_{t-1}$ to model $\phi_{t}$ in eq. (3), and more generally in Def. 2.1, instead of terms such as $Y^{2}_{t-1}$ and $|Y_{t-1}|$ considered for the continuous GARCH models and their modification. This is so because $Y_{t-1}$ is non-negative in our case (count time series).
Remark 2.2.
Some particular cases are obtained from the linear tv-DINGARCH class defined above. For $\alpha_{1}=\alpha_{2}=0$, we obtain the ordinary linear mixed Poisson INGARCH (Christou and Fokianos, 2015; Silva and Barreto-Souza, 2019) models as particular cases. Additionally, by taking $\alpha_{0}\rightarrow 0^{+}$, we also obtain the Poisson INGARCH model as a limiting member of our proposed class. Another interesting and novel model arises when $\beta_{1}=\beta_{2}=0$. Under this setting, the mean of the INGARCH process is constant and the variance is time-dependent as in the case of ordinary GARCH models (Bollerslev, 1986). Such property does not hold for the standard Poisson INGARCH model (1). In this case, we refer to (3) as the Pure INGARCH (P-INGARCH) process. We use the term “pure" to connect the fact that our model mimics the traditional continuous GARCH models (constant mean and time-varying variance). Hence, our approach is general and encompasses many different models studied earlier in the literature.
Some simulated trajectories of the linear tv-DINGARCH models for some parameter settings are shown in the Supplementary Material.
2.1 Stationarity and Ergodicity
[…]
Conditions for the existence and stationarity of the process (3) can be established, for example, by following the strategy by Christou and Fokianos (2014), which relies on establishing weak dependence by Doukhan and Wintenberger (2008). We here obtain such desirable properties based on another approach, the e-chains theory (Meyn and Tweedie, 1993), as follows.
The dynamic latent processes $\{\lambda_{t}\}_{t\geq 1}$ and $\{\phi_{t}\}_{t\geq 1}$ given in (3) can be rewritten in a matrix form. By defining $\boldsymbol{\xi}_{t}=(\lambda_{t},\phi_{t})^{\top}$, for $t\geq 1$, $\boldsymbol{\tau}=(\beta_{0},\alpha_{0})^{\top}$ ${\bf A}=\begin{pmatrix}\beta_{2}&0\\ 0&\alpha_{2}\end{pmatrix}$, and ${\bf B}=\begin{pmatrix}\beta_{1}&0\\ 0&\alpha_{1}\end{pmatrix}$, we have that $\boldsymbol{\xi}_{t}=\boldsymbol{\tau}+{\bf B}(Y_{t-1},Y_{t-1})^{\top}+{\bf A}\boldsymbol{\xi}_{t-1}$. Under this matrix representation, we can adapt the strategy in Liu (2012) to establish the existence and stationarity of our process. In that work, the author provided stochastic properties of a bivariate Poisson INGARCH model. The key point is to show that $\{\boldsymbol{\xi}_{t}\}_{t\in\mathbb{N}}$ is an e-chain Meyn and Tweedie (1993, Ch.18). The following theorem holds true for any mixed Poisson model including the Negative Binomial model.
Theorem 2.1.
Let $\{Y_{t}\}_{t\in\mathbb{N}}$ be a tv-DINGARCH process as in (3). If there exists $p\in[1,\infty]$ such that $\|{\bf A}\|_{p}+2^{1-1/p}\|{\bf B}\|_{p}<1$, then the trivariate process $\{(Y_{t},\lambda_{t},\phi_{t})\}_{t\in\mathbb{N}}$ has a unique stationary and ergodic solution.
[…]
$\lambda_{t}=\beta_{0}+\sum_{i=1}^{p_{1}}\beta_{1i}Y_{t-i}+\sum_{j=1}^{p_{2}}\beta_{2j}\lambda_{t-j},\quad\phi_{t}=\alpha_{0}+\sum_{i=1}^{q_{1}}\alpha_{1i}Y_{t-i}+\sum_{j=1}^{q_{2}}\alpha_{2j}\phi_{t-j}.$ (4)
Then by considering $\boldsymbol{\xi}_{t}=(\lambda_{t},\phi_{t})^{\top}=\boldsymbol{\tau}+\sum_{j=1}^{q}{\bf B}_{j}(Y_{t-j},Y_{t-j})^{\top}+\sum_{i=1}^{m}{\bf A}_{i}\boldsymbol{\xi}_{t-i}$, where $\{{\bf A}_{i}\}_{i=1}^{m}$ and $\{{\bf B}_{j}\}_{j=1}^{q}$ are
${\bf A}_{j}=\begin{pmatrix}\beta_{2j}&0\\ 0&\alpha_{2j}\end{pmatrix}\leavevmode\nobreak\ \mbox{and}\leavevmode\nobreak\ {\bf B}_{j}=\begin{pmatrix}\beta_{1j}&0\\ 0&\alpha_{1j}\end{pmatrix},$
$m=\max(p_{1},p_{2})$, and $q=\max(q_{1},q_{2})$. Then, following the same steps as in the proof of Theorem 2.1 the following result can be established.
Corollary 2.1.
[…]
We close this section by proving finiteness of the moments (see Appendix for a proof) of the first-order negative binomial INGARCH process, which will be used in the next section to establish the asymptotic normality of the conditional maximum likelihood estimators.
Theorem 2.2.
For $\|{\bf A}\|_{1}+\|{\bf B}\|_{1}<1$, the first-order negative binomial INGARCH process $\{Y_{t}\}_{t\in\mathbb{N}}$ given by (3) has finite moments of order $k\in\mathbb{N}$, for an arbitrary $k$.
3 Estimation and Asymptotic Theory
[…]
3.1 Estimators and Asymptotics
Let $\boldsymbol{\theta}=(\theta_{1},\ldots,\theta_{6})^{\top}=(\beta_{0},\beta_{1},\beta_{2},\alpha_{0},\alpha_{1},\alpha_{2})^{\top}$ be the parameter vector, and $y_{1},\ldots,y_{n}$ be a realization of a NB tv-DINGARCH process $\{Y_{t}\}_{t=1}^{n}$. The conditional log-likelihood function of $Y_{2},\ldots,Y_{n}$ given $Y_{1}=y_{1}$ is given by $\ell(\boldsymbol{\theta})=\sum_{t=2}^{n}\ell_{t}(\boldsymbol{\theta})$, where
$\ell_{t}(\boldsymbol{\theta})=y_{t}[\log\lambda_{t}-\log(\lambda_{t}+\phi_{t})]+\phi_{t}[\log\phi_{t}-\log(\lambda_{t}+\phi_{t})]+\log\Gamma(y_{t}+\phi_{t})-\log\Gamma(\phi_{t})-\log y_{t}!,$
for $t=2,\ldots,n$, where we have omitted the dependence of $\lambda_{t}$ and $\phi_{t}$ on $\boldsymbol{\theta}$ for simplicity of notation. In practice, it is necessary to set some initial values for $\lambda_{1}$ and $\phi_{1}$, which are fixed in this paper. More specifically, we get such initial values based on the two first empirical moments of the count time series. This strategy has worked well as demonstrated in the simulated results to be discussed in the sequence. The conditional maximum likelihood estimator (CMLE) of $\boldsymbol{\theta}$, say $\widehat{\boldsymbol{\theta}}$, is given by $\widehat{\boldsymbol{\theta}}=\mbox{argmax}_{\boldsymbol{\theta}\in\Theta}\,\ell(\boldsymbol{\theta})$, where $\Theta=(0,\infty)\times[0,\infty)^{2}\times(0,\infty)\times[0,\infty)^{2}$ denotes the parameter space.
The score function associated to the conditional log-likelihood function is ${\bf U}(\boldsymbol{\theta})=\partial\ell(\boldsymbol{\theta})/\partial\boldsymbol{\theta}=\sum_{t=2}^{n}{\bf U}_{t}(\boldsymbol{\theta})$, with
${\bf U}_{t}(\boldsymbol{\theta})=\left(S_{1t}\dfrac{\partial\lambda_{t}}{\partial\beta_{0}},S_{1t}\dfrac{\partial\lambda_{t}}{\partial\beta_{1}},S_{1t}\dfrac{\partial\lambda_{t}}{\partial\beta_{2}},S_{2t}\dfrac{\partial\phi_{t}}{\partial\alpha_{0}},S_{2t}\dfrac{\partial\phi_{t}}{\partial\alpha_{1}},S_{2t}\dfrac{\partial\phi_{t}}{\partial\alpha_{2}}\right)^{\top},$
and
$S_{1t}=\dfrac{\phi_{t}(y_{t}-\lambda_{t})}{\lambda_{t}(\lambda_{t}+\phi_{t})},\quad S_{2t}=-\dfrac{y_{t}-\lambda_{t}}{\lambda_{t}+\phi_{t}}+\log\left(\dfrac{\phi_{t}}{\lambda_{t}+\phi_{t}}\right)+\Psi(y_{t}+\phi_{t})-\Psi(\phi_{t}),$
for $t=2,\ldots,n$, where $\Psi(x)=d\log\Gamma(x)/dx$, for $x>0$, is the digamma function. Explicit expressions for the derivatives involving $\lambda_{t}$ and $\phi_{t}$ are presented in the Supplementary Material. To establish the asymptotic normality of the CMLEs, the following proposition is necessary.
Proposition 3.1.
Let ${\bf U}_{t}(\boldsymbol{\theta})$ be the $t$-th term of the score function, for $t\geq 2$. For $\|{\bf A}\|_{1}+\|{\bf B}\|_{1}<1$, $E({\bf U}_{t}(\boldsymbol{\theta}))=0$ for all $t\geq 2$, where the expectation is taken regarding the model with true parameter vector $\boldsymbol{\theta}$.
Proof.
Since $E(Y_{t}|\mathcal{F}_{t-1})=\lambda_{t}$ by definition, we immediately obtain that $E(S_{1t}(\boldsymbol{\theta})|\mathcal{F}_{t-1})=0$. We now compute the conditional expectation of $S_{2t}$ given $\mathcal{F}_{t-1}$, which involves $E\left(\Psi(Y_{t}+\phi_{t})|\mathcal{F}_{t-1}\right)$. For $a>0$ and $|c|<1$, we have that
$\sum_{y=0}^{\infty}\dfrac{\Gamma^{\prime}(y+a)}{y!}c^{y}=\dfrac{d}{da}\sum_{y=0}^{\infty}\dfrac{\Gamma(y+a)}{y!}c^{y}=\dfrac{d}{da}\dfrac{\Gamma(a)}{(1-c)^{a}}=\dfrac{\Gamma(a)}{(1-c)^{a}}\left\{\Psi(a)-\log(1-c)\right\},$
where $\Gamma^{\prime}(a)=d\Gamma(a)/da$. Using the above result, it follows that
$E\left(\Psi(Y_{t}+\phi_{t})|\mathcal{F}_{t-1}\right)=\dfrac{(\lambda_{t}\phi_{t}^{-1}+1)^{-\phi_{t}}}{\Gamma(\phi_{t})}\sum_{y=0}^{\infty}\dfrac{\Gamma^{\prime}(y+\phi_{t})}{y!}\left(\dfrac{\lambda_{t}}{\lambda_{t}+\phi_{t}}\right)^{y}=\Psi(\phi_{t})-\log\left(\dfrac{\phi_{t}}{\lambda_{t}+\phi_{t}}\right),$
and then we obtain that $E(S_{2t}(\boldsymbol{\theta})|\mathcal{F}_{t-1})=0$. Hence, for all $t\geq 1$, $E({\bf U}_{t}(\boldsymbol{\theta})|\mathcal{F}_{t-1})=0$.
We will now show that ${\bf U}_{t}(\boldsymbol{\theta})$ is integrable. We have that $|S_{1t}|\leq|y_{t}/\lambda_{t}-1|\leq y_{t}/\beta_{0}+1$. Furthermore, $\partial\lambda_{t}/\partial\beta_{i}$ is a linear combination of $y_{t-1},\ldots,y_{1}$ and $\lambda_{1}$, for $i=0,1,2$. Using these results and under condition $\|{\bf A}\|_{1}+\|{\bf B}\|_{1}<1$, Theorem 2.2 gives that the score function associated with $\beta_{0},\beta_{1}$ and $\beta_{2}$ is integrable. We now study the score function associated with $\alpha_{0}$, $\alpha_{1}$ and $\alpha_{2}$. Note that $\left|-\dfrac{y_{t}-\lambda_{t}}{\lambda_{t}+\phi_{t}}\right|\leq\dfrac{y_{t}}{\lambda_{t}+\phi_{t}}+1\leq\dfrac{y_{t}}{2}\left(\dfrac{1}{\beta_{0}}+\dfrac{1}{\alpha_{0}}\right)+1$, where we have used that fact that $\dfrac{1}{a+b}\leq\dfrac{1}{2}\left(\dfrac{1}{a}+\dfrac{1}{b}\right)$ for $a,b>0$. Also, by using that $\dfrac{1}{2x}<\log x-\Psi(x)<\dfrac{1}{x}$ for $x>0$ (Alzer, 1997), we obtain that $0<\log\phi_{t}-\Psi(\phi_{t})<\phi_{t}^{-1}<\alpha_{0}^{-1}$. Similarly, it follows that $|\Psi(\lambda_{t}+\phi_{t})-\log(\lambda_{t}+\phi_{t})|<(\lambda_{t}+\phi_{t})^{-1}\leq\dfrac{1}{2}\left(\dfrac{1}{\beta_{0}}+\dfrac{1}{\alpha_{0}}\right)$. Therefore, $|S_{2t}|$ is bounded above by a linear combination of $y_{t}$. Moreover, $\dfrac{\partial\phi_{t}}{\partial\alpha_{j}}$ is a linear combination of $y_{t-1},\ldots,y_{1}$ and $\phi_{1}$, for $j=0,1,2$. Again, an application of Theorem 2.2 gives that the score function associated with $\alpha_{0}$, $\alpha_{1}$ and $\alpha_{2}$ is also integrable. ∎
Let us now discuss the asymptotic distribution of the CMLE for the NB tv-DINGARCH model. A more detailed technical result will be established in Theorem 3.1. By using Proposition 3.1 and the square integrability of ${\bf U}(\boldsymbol{\theta})$ (this can be demonstrated using very similar arguments to the proof of Proposition 3.1), we can apply the Central Limit Theorem for martingale difference (for instance, see Corollary 3.1 from Hall and Heyde (1980)) to obtain that
$n^{-1/2}{\bf U}(\boldsymbol{\theta})\stackrel{{\scriptstyle d}}{{\longrightarrow}}N({\bf 0},\boldsymbol{\Omega}_{1}(\boldsymbol{\theta})),\quad\boldsymbol{\Omega}_{1}(\boldsymbol{\theta})\equiv\mbox{plim}_{n\rightarrow\infty}{\bf J}_{1}(\boldsymbol{\theta}),$ (5)
as $n\rightarrow\infty$, with
${\bf J}_{1}(\boldsymbol{\theta})\equiv\dfrac{1}{n}\sum_{t=2}^{n}\mbox{Var}({\bf U}_{t}(\boldsymbol{\theta})|\mathcal{F}_{t-1})=\dfrac{1}{n}\sum_{t=2}^{n}\begin{pmatrix}b_{t}\dfrac{\partial\lambda_{t}}{\partial\boldsymbol{\theta}_{*}}\dfrac{\partial\lambda_{t}}{\partial\boldsymbol{\theta}_{*}}^{\top}&{\bf 0}\\ {\bf 0}&l_{t}\dfrac{\partial\phi_{t}}{\partial\boldsymbol{\theta}^{*}}\dfrac{\partial\phi_{t}}{\partial\boldsymbol{\theta}^{*}}^{\top}\end{pmatrix},$
$\boldsymbol{\theta}_{*}=(\theta_{1},\theta_{2},\theta_{3})^{\top}$, $\boldsymbol{\theta}^{*}=(\theta_{4},\theta_{5},\theta_{6})^{\top}$, $\Psi^{\prime}(x)=d\Psi(x)/dx$,
$b_{t}=\dfrac{\phi_{t}}{\lambda_{t}(\lambda_{t}+\phi_{t})},\quad l_{t}=E\left(\Psi^{2}(Y_{t}+\phi_{t})|\mathcal{F}_{t-1}\right)-\dfrac{\lambda_{t}}{\phi_{t}(\lambda_{t}+\phi_{t})}-\left(\psi(\phi_{t})-\log\left(\dfrac{\phi_{t}}{\lambda_{t}+\phi_{t}}\right)\right)^{2},$
for $t=2,\ldots,n$, where we have used that $E\left(Y_{t}\Psi(Y_{t}+\phi_{t})|\mathcal{F}_{t-1}\right)=\lambda_{t}\left\{\Psi(\phi_{t}+1)-\log\left(\dfrac{\phi_{t}}{\lambda_{t}+\phi_{t}}\right)\right\}$ and that $\Psi(x+1)=\Psi(x)+x^{-1}$, for $x>0$.
Then, we apply the Law of Large Numbers for stationary and ergodic sequences to obtain that
$n^{-1}\dfrac{\partial{\bf U}(\boldsymbol{\theta})}{\partial\boldsymbol{\theta}}\stackrel{{\scriptstyle p}}{{\longrightarrow}}\boldsymbol{\Omega}_{2}(\boldsymbol{\theta}),\quad\boldsymbol{\Omega}_{2}(\boldsymbol{\theta})\equiv\mbox{plim}_{n\rightarrow\infty}{\bf J}_{2}(\boldsymbol{\theta}),$ (6)
as $n\rightarrow\infty$, where
${\bf J}_{2}(\boldsymbol{\theta})\equiv\dfrac{1}{n}\sum_{t=2}^{n}E(-\nabla{\bf U}_{t}(\boldsymbol{\theta})|\mathcal{F}_{t-1})=\dfrac{1}{n}\sum_{t=2}^{n}\begin{pmatrix}b_{t}\dfrac{\partial\lambda_{t}}{\partial\boldsymbol{\theta}_{*}}\dfrac{\partial\lambda_{t}}{\partial\boldsymbol{\theta}_{*}}^{\top}&\bf 0\\ \bf 0&d_{t}\dfrac{\partial\phi_{t}}{\partial\boldsymbol{\theta}^{*}}\dfrac{\partial\phi_{t}}{\partial\boldsymbol{\theta}^{*}}^{\top}\end{pmatrix},$
with
$d_{t}=\Psi^{\prime}(\phi_{t})-E\left(\Psi^{\prime}(Y_{t}+\phi_{t})|\mathcal{F}_{t-1}\right)-\dfrac{\lambda_{t}}{\phi_{t}(\lambda_{t}+\phi_{t})},\quad t=2,\ldots,n,$
where $\Psi^{\prime}(x)=d\Psi(x)/dx$ is the trigamma function. By using the facts that $\Psi^{\prime}(x)=\dfrac{\Gamma^{\prime\prime}(x)}{\Gamma(x)}-\Psi^{2}(x)$ and
$\sum_{y=0}^{\infty}\dfrac{\Gamma^{\prime\prime}(y+a)}{y!}c^{y}=\dfrac{\Gamma(a)}{(1-c)^{a}}\left\{\Psi^{\prime}(a)+[\Psi(a)-\log(1-c)]^{2}\right\},$
for $a>0$ and $|c|<1$, we obtain that $l_{t}=d_{t}$ for all $t$. Therefore, ${\bf J}_{1}(\boldsymbol{\theta})={\bf J}_{2}(\boldsymbol{\theta})$ and $\boldsymbol{\Omega}_{1}(\boldsymbol{\theta})=\boldsymbol{\Omega}_{2}(\boldsymbol{\theta})$.
By combining the above results and using Taylor’s expansion to obtain that
$\sqrt{n}(\widehat{\boldsymbol{\theta}}-\boldsymbol{\theta})\stackrel{{\scriptstyle d}}{{\longrightarrow}}N({\bf 0},\boldsymbol{\Sigma}(\boldsymbol{\theta}))\,\,\mbox{as}\,\,n\rightarrow\infty,$ (7)
with $\boldsymbol{\Sigma}(\boldsymbol{\theta})=\boldsymbol{\Omega}_{2}^{-1}(\boldsymbol{\theta})\boldsymbol{\Omega}_{1}(\boldsymbol{\theta})\boldsymbol{\Omega}_{2}^{-1}(\boldsymbol{\theta})=\boldsymbol{\Omega}_{1}^{-1}(\boldsymbol{\theta})$. We have that $\widehat{\boldsymbol{\Sigma}}={\bf J}_{1}^{-1}(\widehat{\boldsymbol{\theta}})$ is a consistent estimator for $\boldsymbol{\Sigma}$. Another two consistent estimators for $\boldsymbol{\Sigma}$ are ${\bf S}^{-1}_{1}(\widehat{\boldsymbol{\theta}})$ and ${\bf S}^{-1}_{2}(\widehat{\boldsymbol{\theta}})$, where ${\bf S}_{1}(\boldsymbol{\theta})=n^{-1}\sum_{t=2}^{n}{\bf U}_{t}(\boldsymbol{\theta}){\bf U}_{t}(\boldsymbol{\theta})^{\top}$ and ${\bf S}_{2}(\boldsymbol{\theta})=-n^{-1}\sum_{t=2}^{n}\nabla{\bf U}_{t}(\boldsymbol{\theta})$.
There are required additional technical conditions to ensure the consistency and asymptotic normality of the CML estimators, which are provided in the next theorem, where its proof can be found in the Appendix. Before doing that, we have to consider lower and upper bounds for the possible values that the parameters can assume and then we define
=$\mathcal{B}=\{\boldsymbol{\theta}:\beta_{i,low}<\beta_{i}<\beta_{i,up},\,\,\alpha_{j,low}<\alpha_{j}<\alpha_{j,up},\,\,i,j=0,1,2\}.$ (8)
Theorem 3.1.
Let $\{Y_{t}\}_{t=1}^{n}$ follow the linear negative binomial tv-DINGARCH(1,1,1,1) process. Assume that $\|{\bf A}\|_{1}+\|{\bf B}\|_{1}<1$ and that the true value of $\boldsymbol{\theta}$ is an interior point of $\mathcal{B}$. Then, there exists an open set $\mathcal{A}\subset\mathcal{B}$ such that the conditional log-likelihood function $\ell(\cdot)$ has a global maximum point on $\mathcal{A}$, say $\widehat{\boldsymbol{\theta}}$, with probability tending to 1 as $n\rightarrow\infty$. Furthermore, $\widehat{\boldsymbol{\theta}}$ is consistent for $\boldsymbol{\theta}$ and satisfy the asymptotic normality (7).
3.2 Monte Carlo Simulation
[…]
Table 1 summarises the results obtained for settings I and II. Empirical means and standard deviation (SD) of the CMLEs are provided by sample size. TResults show adequate performance of the estimation method, with empirical Monte Carlo means close to the values and standard deviation decreasing with the increase in sample size, as expected.
Setting I $\beta_{0}=15$ $\alpha_{0}=0.5$ $\beta_{1}=0.2$ $\beta_{2}=0.25$ $\alpha_{1}=0.1$ $\alpha_{2}=0.3$
$n=500$ Mean 15.743 0.701 0.196 0.226 0.102 0.254
SD 3.900 0.581 0.045 0.153 0.023 0.153
$n=1000$ Mean 15.491 0.590 0.199 0.233 0.100 0.285
SD 2.973 0.408 0.031 0.115 0.018 0.121
Setting II $\beta_{0}=3$ $\alpha_{0}=0.1$ $\beta_{1}=0.3$ $\beta_{2}=0.15$ $\alpha_{1}=0.2$ $\alpha_{2}=0.3$
$n=500$ Mean 3.126 0.118 0.298 0.130 0.207 0.284
SD 0.515 0.052 0.045 0.092 0.036 0.081
$n=1000$ Mean 3.098 0.107 0.299 0.135 0.203 0.293
SD 0.410 0.034 0.031 0.073 0.024 0.057
Table 1: Empirical mean and standard deviation (SD) of the Monte Carlo estimates for the NB tv-DINGARCH model parameters under Settings I and II. Results are based on 1000 Monte Carlo replications.
Figure 1 illustrates the asymptotic normality of the CMLEs for setting I. Non-parametric density estimator plots of the standardized parameter estimates are displayed alongside the standard Gaussian density (solid line). Dashed and dotted curves indicate densities estimated from the experiments carried out with $n=500$ and $n=1000$, respectively. The density curves of the parameter estimates are mostly overlapping, but some improvement can be seen with the increase of the sample size. Additional density plots for the other parameter configurations can be found in the Supplementary Material.
Figure 1: Non-parametric density plots of standardized parameter estimates due to Setting I under sample sizes $n=500$ and $n=1000$. The solid line corresponds to the standard Gaussian density function.
4 Testing Constant Dispersion
Testing constant value of the dispersion parameter, i.e. $\phi_{t}=\alpha_{0}$ for all $t$ under model (3), is equivalent to testing $H_{0}:\alpha_{1}=\alpha_{2}=0$ against $H_{1}:\alpha_{1}\neq 0\,\,\mbox{or}\,\,\alpha_{2}\neq 0$. Note that the null hypothesis belongs to the boundary of the parameter space, which is a non-standard testing problem; for instance, see Self and Liang (1987), Andrews (2001), and Crainiceanu and Ruppert (2004). We develop and compare heuristically two parametric bootstrap methods; the classical or unrestricted, and the restricted bootstrap recently developed by Cavaliere et al. (2016). The first method considers the usual parametric bootstrap (Efron and Tibshirani, 1993) replications based on the unrestricted CMLEs, while the latter uses the CMLEs under $H_{0}$. Algorithm 1 describes the estimation of the test’s p-value with $B$ replications of the restricted or unrestricted bootstrap.
Input: $\boldsymbol{Y}$ observed count time series data
$B$ bootstrap replications
$\alpha$ significance level
1. Obtain $\widehat{\boldsymbol{\theta}}_{\mathcal{H}_{0}}$ and $\boldsymbol{\widehat{\theta}}$, the model CMLEs under $\mathcal{H}_{0}$ and $\mathcal{H}_{1}$;
2. Compute the observed likelihood ratio $LR=-2(\ell(\widehat{\boldsymbol{\theta}}_{\mathcal{H}_{0}})-\ell(\boldsymbol{\widehat{\theta}}))$;
for $b\leftarrow 1:B$ do
      3A) $\boldsymbol{Y}_{b}\sim\mbox{tv NB-INGARCH}(\widehat{\boldsymbol{\theta}}_{\mathcal{H}_{0}})$;
       // if restricted bootstrap
      3B) $\boldsymbol{Y}_{b}\sim\mbox{tv NB-INGARCH}(\widehat{\boldsymbol{\theta}})$;
       // if unrestricted bootstrap
      4. Obtain $\widehat{\boldsymbol{\theta}}_{\mathcal{H}_{0}}^{b}$ and $\boldsymbol{\widehat{\theta}}^{b}$ fitting tv NB-INGARCH models to $\boldsymbol{Y}_{b}$ under the null and alternative hypothesis;
      5. Let $LR^{b}=-2(\ell(\widehat{\boldsymbol{\theta}}^{b}_{\mathcal{H}_{0}})-\ell(\boldsymbol{\widehat{\theta}}^{b}))$, the replicated LR statistic;
end for
6. If $p_{B}=\sum_{b=1}^{B}I\{LR^{b}>LR\}/B<\alpha$ reject $\mathcal{H}_{0}$;
Algorithm 1 Bootstrap likelihood ratio test of constant ($\mathcal{H}_{0}:\alpha_{1}=\alpha_{2}=0$) versus time-varying dispersion ($\mathcal{H}_{1}:\alpha_{1}\neq 0\mbox{ or }\alpha_{2}\neq 0$) for a tv NB-INGARCH model. Alternatives to step 3 yield restricted (3A) or unrestricted bootstrap (3B) estimators of the test’s p-value, $p_{B}$, where $B$ is the number of replications.
We use a Monte Carlo simulation study to investigate how these methods achieve the desirable significance levels. Time series from the NB tv-DINGARCH process with 200 time observations are simulated under the null hypothesis using four different (varying) settings of the parameter vector $(\beta_{0},\beta_{1},\beta_{2},\alpha_{0})^{\top}$ as follows: (C1) $\beta_{0}=2,\alpha_{0}=1$, $\beta_{1}=0.4,\beta_{2}=0.3$, (C2) $\beta_{0}=2,\alpha_{0}=1$, $\beta_{1}=0,\beta_{2}=0$, (C3) $\beta_{0}=3,\alpha_{0}=0.5$, $\beta_{1}=0.3,\beta_{2}=0.4$, and (C4) $\beta_{0}=3,\alpha_{0}=0.5$, $\beta_{1}=0,\beta_{2}=0$. For each configuration, 500 Monte Carlo replications are used to calculate the empirical significance levels by employing the competing methodologies. The number of replications used to estimate bootstrap $p$-values is $B=500$. Table 2 displays the proportion of times that the restricted and unrestricted bootstrap procedures rejected the null hypothesis. Parameter configurations are set in a way that C2 and C4 are variations of C1 and C3 that do not include effects on the mean. Additional evidence is provided in the Supplementary Material.
Configuration
Significance
level
Restricted
Bootstrap
Unrestricted
Bootstrap
C1: $\beta_{0}=2,\alpha_{0}=1$ $\beta_{1}=0.4,\beta_{2}=0.3$ 0.05 0.046 0.000
0.10 0.088 0.002
C2: $\beta_{0}=2,\alpha_{0}=1$ $\beta_{1}=0,\beta_{2}=0$ 0.05 0.052 0.012
0.10 0.098 0.036
C3: $\beta_{0}=3,\alpha_{0}=0.5$ $\beta_{1}=0.3,\beta_{2}=0.4$ 0.05 0.064 0.002
0.10 0.104 0.002
C4: $\beta_{0}=3,\alpha_{0}=0.5$ $\beta_{1}=0,\beta_{2}=0$ 0.05 0.042 0.022
0.10 0.094 0.058
Table 2: Nominal significance levels produced by the restricted and unrestricted bootstrap hypothesis tests for $H_{0}:\alpha_{1}=\alpha_{2}=0$ for various parameter configurations.
Notably, the restricted parametric bootstrap agrees with the set significance levels, something that occurs under the four parameter configurations investigated in this study, while the unrestricted parametric bootstrap does not provide satisfactory results by underestimating the significance levels. These results show the importance of considering a restricted bootstrap for testing constant dispersion in the tv-DINGARCH models.
5 Measles Data Analysis
The linear negative binomial tv-DINGARCH model is now applied to modeling the weekly number of reported measles infections in North Rhine-Westphalia, Germany. The series is observed between January 2001 and May 2013 (646 observations), publicly available in the R package tscount. Figure 2 displays $\{Y_{t}\}_{t=1}^{646}$ on the left, and the series autocorrelation function on the right.
Figure 2: On the left, weekly cases of measles reported in North Rhine-Westphalia, Germany, between January 2001 and May 2013. The autocorrelation function of the series is shown on the right.
Fitting of the NB tv-DINGARCH model will be compared to the fit of ordinary NBINGARCH model with an identical mean time series structure. The purpose of this exercise is to assess how a time-varying dispersion changes the model adequacy to this real data. To this end, we will evaluate goodness-of-fit and predictions for each model. $\mathcal{M}_{tv}$ will denote the NB tv-DINGARCH model while $\mathcal{M}_{ord}$ denotes the ordinary case (which assumes a constant dispersion).
[…]
Parameter tv-DINGARCH INGARCH
Estimate (se) 95% CI Estimate (se) 95% CI
$\beta_{0}$ 0.259 (0.048) (0.164, 0.353) 0.194 (0.127) (0.000, 0.444)
$\beta_{1}$ 0.579 (0.034) (0.512, 0.645) 0.583 (0.090) (0.407, 0.759)
$\beta_{2}$ 0.342 (0.039) (0.265, 0.419) 0.390 (0.094) (0.205, 0.574)
$\alpha_{0}$ 0.775 (0.057) (0.663, 0.886) 0.736 (0.110) (0.521, 0.952)
$\alpha_{1}$ 0.079 (0.007) (0.065, 0.093)
$\alpha_{2}$ 0.000 (0.010) (0.000, 0.020)
Table 3: Conditional maximum likelihood estimates, standard errors (se), and 95% confidence intervals for log-linear tv-DINGARCH and ordinary log-linear INGARCH models applied to weekly counts of measles in Germany.
In addition, we consider Probability Integral Transform (PIT) plots (Czado et al., 2009), an approach that enables the comparison of count data models via their predictive distributions. A model providing a good fit to the data in this aspect will render a PIT plot resembling a uniform distribution, where major deviations typically indicate problems of overdispersion or underdispersion of the model’s predictive distribution. These are reported in Figure 3 for both models, and while that of $\mathcal{M}_{tv}$ is near the uniform as desired, the upside-down U shaped PIT of $\mathcal{M}_{ord}$ points to a non adequate fit.
Figure 3: Probability Integral Transform (PIT) assessment for $\mathcal{M}_{ord}$ (on the left) and $\mathcal{M}_{tv}$ (to the right).
It is often of interest to practitioners to select between models according to their forecasting power. To that end, we consider a forecasting exercise that explroes one-step-ahead (OSA) prediction from $\mathcal{M}_{tv}$ and $\mathcal{M}_{ord}$. This is done in a recursive manner in order to provide multiple OSA predictions from each model. We start by defining the initial training data from the beginning of the study until October 2004 (week 4) which contains 200 observations. Both models are fitted to the training set and used to predict the next week’s counts (week 1 of November 2004) via the conditional median and mode of the distributions. The conditional (or predictive) distributions are Negative-Binomial($\hat{\mu}_{t+1},\hat{\phi}_{t+1}$), or Negative-Binomial($\hat{\mu}_{t+1},\hat{\phi}$) respectively for $\mathcal{M}_{tv}$ and $\mathcal{M}_{ord}$ and $t=200$ at this step. Once the prediction is obtained, we add week 1 of November 2004 to the training set, refit both models and gather the new OSA predictions. Proceeding until the end of the study period gives a total of 446 predictions from each model. Pseudocode describing the steps to this prediction exercise is given in Algorithm 2.
Input: $\boldsymbol{Y}$ observed count trajectory of length $n$
$n_{0}$ starting point of prediction exercise
0. $s\leftarrow n_{0}$;
while $s<n$ do
      1. $\boldsymbol{Y}(s)\leftarrow Y_{[1:s]}$;
       // train data
      2. Fit the tv-NBINGARCH model to $\boldsymbol{Y}(s)$;
      3. From 2, gather the CMLEs $\widehat{\boldsymbol{\theta}}(s)$ and the fitted $\widehat{\boldsymbol{\lambda}}(s)$ and $\widehat{\boldsymbol{\phi}}(s)$ of step $s$;
      4. Obtain the OSA mean $\widehat{\mu}_{s+1}=\widehat{\beta}_{0}^{s}+\widehat{\beta}_{1}^{s}Y_{s}+\widehat{\beta}_{2}^{s}\widehat{\mu}_{s}$ and $\widehat{\phi}_{s+1}=\widehat{\alpha}_{0}^{s}+\widehat{\alpha}_{1}^{s}Y_{s}+\widehat{\alpha}_{2}^{s}\widehat{\phi}_{s}$;
      5A) $\widehat{Y}_{s+1}=\widehat{\mu}_{s+1}$
       // prediction via mean
      5B) $\widehat{Y}_{s+1}=\texttt{qbinom}(0.5,\texttt{size}=\widehat{\phi}_{s+1},\texttt{mu}=\widehat{\mu}_{s+1})$
       // prediction via median
       $s=s+1$;
       // increment step
end while
Output: $\boldsymbol{\widehat{Y}}$ vector of $(n-n_{0})$ OSA predictions.
Algorithm 2 Recursive algorithm for obtaining $\boldsymbol{\widehat{Y}}$, the one-step-ahead (OSA) predicted values of $\boldsymbol{Y}_{[n_{0}+1:n]}$. The training data at iteration $s$, $\boldsymbol{Y}(s)$, is incremented and the model is refitted to obtain the OSA forecast $\widehat{Y}_{s+1}$. Steps 5A) and 5B) provide alternative prediction methods via the mean or median.
Algorithm 2 is presented in terms of the $\mathcal{M}_{tv}$ model but works similarly for $\mathcal{M}_{ord}$. In this case, $\mathcal{M}_{ord}$ is fitted in step two and the predictions in 5A) and 5B) take in the (fixed) dispersion parameter estimated in step 2. By employing the Algorithm with $\mathcal{M}_{tv}$ and $\mathcal{M}_{ord}$, their predictions are summarized via the root mean forecasting error (RMSFE). Let $n_{0}$ denote the time point chosen to start the prediction exercise; in our case $n_{0}=200$ the RMSFE of the forecasting step $t$ is
$\mbox{RMSFE}_{t}=\sqrt{\frac{1}{t-n_{0}}\sum_{s=n_{0}+1}^{t}(Y_{s}-\widehat{Y}_{s})^{2}},$
where $Y_{s}$ and $\widehat{Y}_{s}$, are the observed and predicted counts, respectively. Calculation of RMSFE, from the 446 total predictions from $\mathcal{M}_{tv}$ and $\mathcal{M}_{ord}$, yields Figure 4. On the left, $\widehat{Y}_{s}$ is the mode of the conditional distribution and on the right it is taken as the median. Prediction by the median yields a smaller prediction error in comparison to the mode, and in both cases the tv-DINGARCH model produces the smallest RMSFE values.
Figure 4: RMSFE of predictions obtained with the fit of ordinary (dashed lines) and time-varying dispersion (solid lines) INGARCH models to the weekly count of measles in North Rhine-Westphalia, Germany. On the left, predicted values are the mode of the predictive distributions, whereas the median is taken as the point on the right.
This empirical illustration on the weekly number of measles cases in North Rhine-Westphalia, Germany, demonstrated that the time-varying dispersion is a promising and important extension of the ordinary INGARCH processes that achieved improvement in terms of goodness-of-fit and forecasting.
6 Conclusions
We proposed a class of time-varying dispersion INGARCH (tv-DINGARCH) models and explore stochastic properties such as stationarity and ergodicity. Estimation of parameters was addressed through conditional maximum likelihood estimation (CMLE) and its associated asymptotic theory was established. Monte Carlo simulations were conducted to evaluate the performance of the CMLE. Moreover, we developed bootstrap methodologies to test for constant dispersion and showed via simulated studies that the restricted bootstrap is preferred over the unrestricted parametric one. We analyzed the weekly number of reported measles cases in north Rhine-Westphalia, Germany, from January 2001 to May 2013, and found that the tv-DINGARCH approach delivers much better results regarding goodness-of-fit and prediction when compared to the ordinary INGARCH model.
A log-linear version of our model allowing for the inclusion of covariates deserves future research. A first-order log-linear tv-DINGARCH process $\{Y_{t}\}_{t\geq 1}$ allowing for covariates/exogenous time series is given as in Definition 2.1 with $\lambda_{t}\equiv\exp(\mu_{t})$ and $\phi_{t}\equiv\exp(\nu_{t})$, where
$\mu_{t}=\beta_{0}+\beta_{1}\log(Y_{t-1}+1)+\beta_{2}\mu_{t-1}+\boldsymbol{\delta}^{\top}{\bf X}_{t},\quad\nu_{t}=\alpha_{0}+\alpha_{1}\log(Y_{t-1}+1)+\alpha_{2}\nu_{t-1}+\boldsymbol{\gamma}^{\top}{\bf W}_{t},$
$\beta_{i},\alpha_{i}\in\mathbb{R}$ for $i=0,1,2$, with ${\bf X}_{t}$ and ${\bf W}_{t}$ covariates with associated real-valued coefficients $\boldsymbol{\delta}$ and $\boldsymbol{\gamma}$. The ordinary log-linear INGARCH model (Fokianos and Tjøstheim, 2011) is obtained as a particular case from the log-linear tv-DINGARCH process by taking $\alpha_{1}=\alpha_{2}=0$ and $\boldsymbol{\gamma}=0$. This topic will be investigated in a future communication. Finally, it is worth mentioning that other mixed Poisson distributions rather than negative binomial can be used for our tv-DINGARCH formulation such as Poisson-inverse Gaussian distribution.
Appendix
A-1 Proof of Theorem 2.1
[…]
Let $w$ be a continuous function with compact support on $(0,\infty)^{2}$ and assumed to be bounded $|w|<1$ without loss of generality. Consider $k=1$, ${\bf x},{\bf z}\in(0,\infty)^{2}$, and $\epsilon>0$ arbitrary. Denote by $f^{mp}_{\boldsymbol{\xi}}(\cdot)$ the probability function of a mixed Poisson distribution with mean $\lambda$ and variance $\lambda+\lambda^{2}/\phi$, where $\boldsymbol{\xi}=(\lambda,\phi)^{\top}$. It follows that
$\left|E(w(\boldsymbol{\xi}_{1})|\boldsymbol{\xi}_{0}={\bf x})-E(w(\boldsymbol{\xi}_{1})|\boldsymbol{\xi}_{0}={\bf z})\right|\leq$
$\sum_{y=0}^{\infty}|w(\boldsymbol{\tau}+{\bf B}(y,y)^{\top}+{\bf A}\boldsymbol{\bf x})f^{mp}_{\bf x}(y)-w(\boldsymbol{\tau}+{\bf B}(y,y)^{\top}+{\bf A}\boldsymbol{\bf z})f^{mp}_{\bf z}(y)|\leq$
$\sum_{y=0}^{\infty}f^{mp}_{\bf x}(y)|w(\boldsymbol{\tau}+{\bf B}(y,y)^{\top}+{\bf A}\boldsymbol{\bf x})-w(\boldsymbol{\tau}+{\bf B}(y,y)^{\top}+{\bf A}\boldsymbol{\bf z})|+$ (A-1)
$\sum_{y=0}^{\infty}|w(\boldsymbol{\tau}+{\bf B}(y,y)^{\top}+{\bf A}\boldsymbol{\bf z})|\,|f^{mp}_{\bf x}(y)-f^{mp}_{\bf z}(y)|.$ (A-2)
We now find an upper bound for the term (A-2). Denote by $f^{pois}_{\lambda}(\cdot)$ the probability function of a Poisson distribution with mean $\lambda$. From the mixed Poisson stochastic representation, we have that $f^{mp}_{\bf x}(y)=E(f^{pois}_{x_{1}Z_{1}}(y))$, where $Z_{1}$ is the associated latent random variables with distribution depending on $x_{2}$. Similar representation holds for $f^{mp}_{\bf z}(y)$ in terms of an associated latent factor $Z_{2}$ with distribution depending on $z_{2}$. Using this and the recalling the fact that $|g|<1$, we obtain that
$\sum_{y=0}^{\infty}|w(\boldsymbol{\tau}+{\bf B}(y,y)^{\top}+{\bf A}\boldsymbol{\bf z})|\,|f^{mp}_{\bf x}(y)-f^{mp}_{\bf z}(y)|\leq\sum_{y=0}^{\infty}|f^{mp}_{\bf x}(y)-f^{mp}_{\bf z}(y)|=$
$\sum_{y=0}^{\infty}|E(f^{pois}_{x_{1}Z_{1}}(y))-E(f^{pois}_{z_{1}Z_{2}}(y))|\leq E\left(\sum_{y=0}^{\infty}|f^{pois}_{x_{1}Z_{1}}(y)-f^{pois}_{x_{2}Z_{2}}(y)|\right).$ (A-3)
By using inequality (A.1) from Wang et al. (2014) (see also Liu (2012)), we have that $\sum_{y=0}^{\infty}|f^{pois}_{x_{1}Z_{1}}(y)-f^{pois}_{x_{2}Z_{2}}(y)|\leq 2\left(1-\exp\{-|x_{1}Z_{1}-x_{2}Z_{2}|\}\right)$. Hence, (A-1) is bounded above by
$E\left(\sum_{y=0}^{\infty}|f^{pois}_{x_{1}Z_{1}}(y)-f^{pois}_{x_{2}Z_{2}}(y)|\right)\leq 2E\left(1-e^{-|x_{1}Z_{1}-x_{2}Z_{2}|}\right)\leq$
$2E\left(1-e^{-|x_{1}-x_{2}|(Z_{1}+Z_{2})}\right)\leq 2E\left(|x_{1}-x_{2}|(Z_{1}+Z_{2})\right)=2|x_{1}-x_{2}|\leq 2\|{\bf x}-{\bf z}\|_{p},$
for $p\in[1,\infty]$, where we have used in the second inequality the fact that $|ab-cd|\leq|a-c|(b+d)$ for $a,b,c,d>0$. In the third inequality, we used that $1-e^{-x}\leq x$ for all $x>0$. Finally, the fourth inequality follows, for instance, by (Liu, 2012, pp. 108).
The upper bound for the term given in (A-1) follows exactly as discussed in Liu (2012): choose $\epsilon^{\prime}>0$ and $\eta>0$ small enough such as $\epsilon^{\prime}+\dfrac{8\eta}{1-\|{\bf A}\|_{p}}<\epsilon$ and $\|{\bf x}-{\bf z}\|_{p}<\eta$ implying $|w({\bf x})-w({\bf z})|<\epsilon^{\prime}$, with $p\in[1,\infty]$. In this part, we are using the fact that $\|{\bf A}\|_{p}<1$ for some $p\in[1,\infty]$, which follows from the assumption that there exists $p$ such that $\|{\bf A}\|_{p}+2^{1-1/p}\|{\bf B}\|_{p}<1$.
By combining the above results, we obtain that $\left|E(w(\boldsymbol{\xi}_{1})|\boldsymbol{\xi}_{0}={\bf x})-E(w(\boldsymbol{\xi}_{1})|\boldsymbol{\xi}_{0}={\bf z})\right|<\epsilon^{\prime}+2\|{\bf x}-{\bf z}\|_{p}$. For general $k\geq 2$, the result follows by using mathematical induction, exactly as done in Chapter 4 from Liu (2012), and therefore it is omitted. With the e-chain property established for the bivariate process $\{\boldsymbol{\xi}_{t}\}$ and under the assumption of existence of $p\in[1,\infty]$ such that $\|{\bf A}\|_{p}+2^{1-1/p}\|{\bf B}\|_{p}<1$, following the same steps that proofs of Proposition 4.2.1(b) and Proposition 4.3.1 from Liu (2012), we obtain the desired result.
A-2 Proof of Theorem 2.2
Let $\{Y_{t}\}_{t\in\mathbb{N}}$ be a first-order negative binomial INGARCH process. Under condition $\|{\bf A}\|_{1}+\|{\bf B}\|_{1}<1$, Theorem 2.1 holds (for $p=1$) and gives us that $\{Y_{t}\}_{t\in\mathbb{N}}$ has a unique stationary and ergodic solution. Moreover, this process is a Markov chain. Therefore, we will now show that Tweedie’s criterion is satisfied for a test function to show the finiteness of arbitrary moments stated in the lemma (Meyn and Tweedie, 1993).
Consider the test function $V(x,z)=1+x^{k}+z^{k}$, for $x,z>0$ and arbitrary $k\in\mathbb{N}$. It follows that
$E\left(V(\lambda_{t},\phi_{t})|\lambda_{t-1}=\lambda,\phi_{t-1}=\phi\right)$ $=$ $1+E\left(\lambda_{t}^{k}|\lambda_{t-1}=\lambda,\phi_{t-1}=\phi\right)+E\left(\phi_{t}^{k}|\lambda_{t-1}=\lambda,\phi_{t-1}=\phi\right)$ (A-4)
$=$ $1+E\left((\beta_{0}+\beta_{1}Y+\beta_{2}\lambda)^{k}\right)+E\left((\alpha_{0}+\alpha_{1}Y+\alpha_{2}\phi)^{k}\right),$
where $Y\sim\mbox{NB}(\lambda,\phi)$. From proof of Lemma (A.1) from Christou and Fokianos (2014), we have that the $d$-th moment of $Y\sim\mbox{NB}(\lambda,\phi)$, for $d\geq 2$, satisfies the recursive equation
$E(Y^{d})=\lambda\left\{1+\sum_{j=1}^{d-1}\left[\binom{d-1}{j}+\dfrac{1}{\phi}\binom{d-1}{j-1}\right]E(Y^{j})\right\},$ (A-5)
with $E(Y)=\lambda$.
From (A-5), we immediately obtain that $E(Y^{d})=\lambda^{d}(1+1/\phi)+O(\lambda^{d-1})$ as $\lambda\rightarrow\infty$. Using this result, it follows that
$E\left((\beta_{0}+\beta_{1}Y+\beta_{2}\lambda)^{k}\right)$ $=$ $\sum_{j=0}^{k}\binom{k}{j}(\beta_{0}+\beta_{2}\lambda)^{k-j}\beta_{1}^{j}E(Y^{j})=\sum_{j=0}^{k}\binom{k}{j}(\beta_{2}\lambda)^{k-j}\beta_{1}^{j}\lambda^{j}+O(\lambda^{k-1})$
$=$ $\lambda^{k}(\beta_{1}+\beta_{2})^{k}+O(\lambda^{k-1}).$
Similarly, we obtain that
$E\left((\alpha_{0}+\alpha_{1}Y+\alpha_{2}\phi)^{k}\right)=(\alpha_{1}\lambda+\alpha_{2}\phi)^{k}+O(\lambda^{k-1}).$
Hence, we can express the conditional expectation (A-4) as
$E\left(V(\lambda_{t},\phi_{t})|\lambda_{t-1}=\lambda,\phi_{t-1}=\phi\right)=V(\lambda,\phi)\dfrac{1+\lambda^{k}(\beta_{1}+\beta_{2})^{k}+(\alpha_{1}\lambda+\alpha_{2}\phi)^{k}+O(\lambda^{k-1})}{1+\lambda^{k}+\phi^{k}}.$ (A-6)
We will now analyze the behavior of the terms multiplying $V(\lambda,\phi)$ in (A-6) for large values of $\lambda$ and $\phi$. First, note that $\lambda^{*}=\dfrac{\beta_{0}}{1-\beta_{2}}$ and $\phi^{*}=\dfrac{\alpha_{0}}{1-\alpha_{2}}$ are the smallest reachable points of $\lambda_{t}$ and $\phi_{t}$, respectively, for all $t$. Hence, we consider $\lambda\in[\lambda^{*},\infty)$ and $\phi\in[\phi^{*},\infty)$, both bounded away from zero. Consider the polar coordinates transformation $\lambda=r\cos\delta$ and $\phi=r\sin\delta$, for $r>0$ and $\delta\in(0,\pi/2)$ such that $\lambda\geq\lambda^{*}$ and $\phi\geq\phi^{*}$. In terms of polar coordinates, we have that
$\dfrac{1+\lambda^{k}(\beta_{1}+\beta_{2})^{k}+(\alpha_{1}\lambda+\alpha_{2}\phi)^{k}}{1+\lambda^{k}+\phi^{k}}$ $=$ $\dfrac{1+r^{k}[(\beta_{1}+\beta_{2})^{k}\cos^{k}\delta+(\alpha_{1}\cos\delta+\alpha_{2}\sin\delta)^{k}]}{1+r^{k}(\cos^{k}\delta+\sin^{k}\delta)}$
$\longrightarrow$ $\dfrac{(\beta_{1}+\beta_{2})^{k}\cos^{k}\delta+(\alpha_{1}\cos\delta+\alpha_{2}\sin\delta)^{k}}{\cos^{k}\delta+\sin^{k}\delta}$
$\leq$ $\dfrac{(\beta_{1}+\beta_{2})^{k}\cos^{k}\delta+\max\{\cos\delta,\sin\delta\}(\alpha_{1}+\alpha_{2})^{k}}{\cos^{k}\delta+\sin^{k}\delta}$
$\leq$ $(\beta_{1}+\beta_{2})^{k}+(\alpha_{1}+\alpha_{2})^{k}$
$\leq$ $\beta_{1}+\beta_{2}+\alpha_{1}+\alpha_{2}$
$=$ $\|{\bf A}\|_{1}+\|{\bf B}\|_{1}<1,$
as $r\rightarrow\infty$, where the last inequality follows from lemma’s assumption implying that $\beta_{1}+\beta_{2}<1$ and $\alpha_{1}+\alpha_{2}<1$. Further, the remaining term converges to 0 as $r\rightarrow\infty$: $\dfrac{O(\lambda^{k-1})}{1+\lambda^{k}+\phi^{k}}=O(r^{k-1})/O(r^{k})=O(r^{-1})$.
Therefore, there exist real constants $\kappa_{1}\in(0,1)$, $\kappa_{2}>0$, and $L>0$ such that (A-4) is bounded above as follows:
$E\left(V(\lambda_{t},\phi_{t})|\lambda_{t-1}=\lambda,\phi_{t-1}=\phi\right)\leq(1-\kappa_{1})V(\lambda,\phi)+\kappa_{2}I\{(\lambda,\phi)\in G\},$
[…]
In other words, Tweedie’s criterion is satisfied and we conclude that the $k$-th moment of $\lambda_{t}$ and $\phi_{t}$ are finite for arbitrary $k\in\mathbb{N}$. Now, using the fact that $Y_{t}|\mathcal{F}_{t-1}\sim\mbox{NB}(\lambda_{t},\phi_{t})$ and Eq. (A-5), we have that the $k$-th conditional moment of $Y_{t}$ given $\mathcal{F}_{t-1}$ is a polynomial on $\lambda_{t}$ and $1/\phi_{t}$; note that $1/\phi_{t}\leq 1/\alpha_{0}$ for all $t$. Since the $l$-th moment of $\lambda_{t}$ is finite for arbitrary $l\in\mathbb{N}$, we obtain that the unconditional $k$-th moment of $Y_{t}$ is finite for all $k\geq 1$ as well.
Lemma A-1.
Let $\mathcal{B}$ as defined in (8). Under the linear negative binomial tv-DINGARCH(1,1,1,1) model with $\|{\bf A}\|_{1}+\|{\bf B}\|_{1}<1$, there exists a sequence of random variables $\{M_{n}\}_{n\in\mathbb{N}}$ such that
$\max_{i,j,k=1,2,\ldots,6}\sup_{\boldsymbol{\theta}\in\mathcal{B}}\bigg{|}\dfrac{1}{n}\dfrac{\partial^{3}\ell(\boldsymbol{\theta})}{\partial\theta_{i}\partial\theta_{j}\partial\theta_{k}}\bigg{|}\leq M_{n},$
where $M_{n}\stackrel{{\scriptstyle p}}{{\longrightarrow}}m$ as $n\rightarrow\infty$, with $m<\infty$ and $\boldsymbol{\theta}=(\theta_{1},\theta_{2},\theta_{3},\theta_{4},\theta_{5},\theta_{6})^{\top}=(\beta_{0},\beta_{1},\beta_{2},\alpha_{0},\alpha_{1},\alpha_{2})^{\top}$.
Proof.
We will show that the result holds for the third derivative of $\ell(\cdot)$ with respect to $\beta_{1}$ and also $\alpha_{1}$. The remaining cases follow the very same steps as these two and therefore are omitted.
The third derivative of $\ell_{t}(\boldsymbol{\theta})$ with respect to $\beta_{1}$ is given by $\dfrac{\partial^{3}\ell_{t}(\boldsymbol{\theta})}{\partial\beta_{1}^{3}}=\dfrac{\partial^{2}S_{1t}}{\partial\phi_{t}^{2}}\left(\dfrac{\partial\lambda_{t}}{\partial\beta_{1}}\right)^{3}$, where
$\dfrac{\partial^{2}S_{1t}}{\partial\phi_{t}^{2}}=\dfrac{2\phi_{t}}{\lambda_{t}^{2}(\lambda_{t}+\phi_{t})^{2}}\left\{\dfrac{[y_{t}(2\lambda_{t}+\phi_{t})-\lambda_{t}^{2}](2\lambda_{t}+\phi_{t})}{\lambda_{t}(\lambda_{t}+\phi_{t})}-(y_{t}-\lambda_{t})\right\}$
and
$\dfrac{\partial\lambda_{t}}{\partial\beta_{1}}=\sum_{j=1}^{t-1}\beta_{2}^{j-1}y_{t-j}.$
It is straightforward to obtain that there are constants $c_{1},c_{2}>0$ that do not depend on $\boldsymbol{\theta}$ such that $\left|\dfrac{\partial^{3}\ell_{t}(\boldsymbol{\theta})}{\partial\beta_{1}^{3}}\right|\leq c_{2}y_{t}+c_{1}$. Also, $\left|\dfrac{\partial\lambda_{t}}{\partial\beta_{1}}\right|\leq\displaystyle\sum_{j=1}^{t-1}\beta_{2,up}^{j-1}y_{t-j}\equiv\mu_{1t}$. Therefore, $\left|\dfrac{\partial^{3}\ell_{t}(\boldsymbol{\theta})}{\partial\beta_{1}^{3}}\right|\leq c_{2}y_{t}\mu_{1t}^{3}+c_{1}\mu_{1t}^{3}\equiv M_{1t}$, where $M_{1t}$ does not depend on $\boldsymbol{\theta}$. Under the assumption $\|{\bf A}\|_{1}+\|{\bf B}\|_{1}<1$ and using Theorem 2.2, the Law of Large Numbers for stationary and ergodic sequences can be applied to obtain that $n^{-1}\displaystyle\sum_{t=2}^{n}M_{1t}\stackrel{{\scriptstyle p}}{{\longrightarrow}}m_{1}$ as $n\rightarrow\infty$, where $m_{1}$ is a finite constant, as similarly argued by Fokianos et al. (2009) (proof of Lemma 3.4 of their Supplementary Material).
The case for $\alpha_{1}$ is more involving. We have that $\dfrac{\partial^{3}\ell_{t}(\boldsymbol{\theta})}{\partial\alpha_{1}^{3}}=\dfrac{\partial^{2}S_{2,t}}{\partial\phi_{t}^{2}}\left(\dfrac{\partial\phi_{t}}{\partial\alpha_{1}}\right)^{3}$, where
$\dfrac{\partial^{2}S_{2,t}}{\partial\phi_{t}^{2}}$ $=$ $-2\dfrac{y_{t}-\lambda_{t}}{(\lambda_{t}+\phi_{t})^{3}}+\dfrac{\lambda_{t}^{2}+2\lambda_{t}\phi_{t}}{\phi_{t}^{2}(\lambda_{t}+\phi_{t})^{2}}+\Psi^{\prime\prime}(y_{t}+\phi_{t})-\Psi^{\prime\prime}(\phi_{t})$ (A-7)
and
$\dfrac{\partial\phi_{t}}{\partial\alpha_{1}}=\displaystyle\sum_{j=1}^{t-1}\alpha_{2}^{j-1}y_{t-j}.$
Inequality (15) from Guo and Qi (2013) gives that $|\Psi^{\prime\prime}(x)|<\dfrac{1}{x}+\dfrac{2}{x^{3}}$, for $x>0$. Hence, it follows that $|\Psi^{\prime\prime}(y_{t}+\phi_{t})|<\dfrac{1}{(y_{t}+\phi_{t})^{2}}+\dfrac{2}{(y_{t}+\phi_{t})^{3}}\leq\dfrac{1}{\phi_{t}^{2}}+\dfrac{2}{\phi_{t}^{3}}\leq\dfrac{1}{\alpha_{0}^{2}}+\dfrac{2}{\alpha_{0}^{3}}\leq\dfrac{1}{\alpha_{0,inf}^{2}}+\dfrac{2}{\alpha_{0,inf}^{3}}$. Similarly, $|\Psi^{\prime\prime}(\phi_{t})|<\dfrac{1}{\alpha_{0,inf}^{2}}+\dfrac{2}{\alpha_{0,inf}^{3}}$. The modulus of the remaining terms in (A-7) are bounded above (as done for the third derivative involving $\beta_{1}$) by a linear function of $y_{t}$ with positive coefficients do not depend on $\boldsymbol{\theta}$. Moreover, $\left|\dfrac{\partial\phi_{t}}{\partial\alpha_{1}}\right|\leq\displaystyle\sum_{j=1}^{t-1}\alpha_{2,up}^{j-1}y_{t-j}\equiv\mu_{2t}$.
From the above results, we conclude that there exist constants $c_{3}>0$ and $c_{4}>0$ that do not depend on $\boldsymbol{\theta}$ such that $\left|\dfrac{\partial^{3}\ell_{t}(\boldsymbol{\theta})}{\partial\alpha_{1}^{3}}\right|\leq c_{4}y_{t}\mu_{2t}^{3}+c_{3}\mu_{2t}^{3}\equiv M_{2t}$. Now, by arguing exactly as done for the case $\beta_{1}$ above, we obtain that $n^{-1}\sum_{t=2}^{n}M_{2t}\stackrel{{\scriptstyle p}}{{\longrightarrow}}m_{2}$ when $n\rightarrow\infty$, where $m_{2}$ is a finite constant. ∎
A-3 Proof of Theorem 3.1
We will show that the Conditions (A1), (A2), and (A3) from Lemma 3.1 by Jensen and Rahbek (2004) are satisfied, which give us the desired results stated in the theorem. As argued before for the general case, we use Proposition 3.1 and the Central Limit Theorem for martingale difference to establish the weak convergence involving the score function in (5). The existence and finiteness of the matrix $\boldsymbol{\Omega}_{1}(\boldsymbol{\theta})$ follow from Theorem 2.2 that provides the finiteness of the moments of all orders for the linear NB tv-DINARCH process under the assumption that $\|{\bf A}\|_{1}+\|{\bf B}\|_{1}<1$, which is in force. Therefore, Condition (A1) holds.
Moreover, under the assumption $\|{\bf A}\|_{1}+\|{\bf B}\|_{1}<1$, we obtain from Theorem 2.1 that $\{Y_{t}\}_{t\in\mathbb{N}}$ is a stationary and ergodic sequence. Then, the Law of Large Numbers for stationary and ergodic sequences gives us that (6) is valid, with the existence and finiteness of the matrix $\boldsymbol{\Omega}_{2}(\boldsymbol{\theta})$ being ensured again by Theorem 2.2. That is, Condition (A2) from Lemma 3.1 by Jensen and Rahbek (2004) is satisfied. Finally, Condition (A3) has been established in Lemma A-1.
Acknowledgments
H. Ombao would like to acknowledge support by KAUST Research Fund. L.S.C. Piancastelli wishes to acknowledge the financial support of Science Foundation Ireland.

### v3

@@
 Abstract
-We propose a general class of INteger-valued Generalized AutoRegressive Conditionally Heteroscedastic (INGARCH) processes by allowing time-varying mean and dispersion parameters, which we call time-varying dispersion INGARCH (tv-DINGARCH) models.
-More specifically, we consider mixed Poisson INGARCH models and allow for dynamic modeling of the dispersion parameter (as well as the mean), similar to the spirit of the ordinary GARCH models.
-We derive conditions to obtain first and second-order stationarity, and ergodicity as well.
-Estimation of the parameters is addressed and their associated asymptotic properties are established as well.
+We introduce a general class of INteger-valued Generalized AutoRegressive Conditionally Heteroscedastic (INGARCH) processes by allowing simultaneously time-varying mean and dispersion parameters.
+We call such models time-varying dispersion INGARCH (tv-DINGARCH) models.
+More specifically, we consider mixed Poisson INGARCH models and allow for dynamic modeling of both mean and dispersion parameters.
+We derive conditions to obtain stochastic stability of tv-DINGARCH processes.
+Additionally, we study maximum likelihood estimation in detail including its asymptotic distribution.
 A restricted bootstrap procedure is proposed for testing constant dispersion against time-varying dispersion.
@@
 Keywords: Autocorrelation; Count time series; Overdispersion; Time-varying dispersion parameter; Volatility.
+MOS subject classifications: 62M09; 62M10.
 1 Introduction
-Modeling count time series data is a challenging and very exciting research topic with applications in many different areas such as epidemiology, sociology, economics, and health science.
-A well-established methodology, for dealing with count time series data, has been developed under the framework of INteger-valued Generalized AutoRegressive Conditional Heteroscedastic (INGARCH) models, which have been initially studied and explored by Heinen (2003), Ferland et al.
+Modeling count time series data is a challenging and exciting research topic with applications in many areas such as epidemiology, sociology, economics, and health science.
+A well-established methodology, for dealing with count time series data, has been developed under the framework of the so-called INteger-valued Generalized AutoRegressive Conditional Heteroscedastic (INGARCH) models, which have been initially studied and explored by Heinen (2003), Ferland et al.
 (2006), Fokianos et al.
 (2009), and Fokianos (2011).
-The INGARCH nomenclature emerges from the fact that for a Poisson distribution (this assumption is imposed in the aforementioned papers), the mean equals its variance, therefore modeling of mean implies modeling of variance like the classic continuous GARCH models introduced by Bollerslev (1986).
-Consequently, such models are considered in the literature as integer counterpart of the GARCH models although this terminology should be used cautiously because ordinary GARCH processes do not consider dynamics for the conditional mean; for instance, see Fokianos et al.
-(2009).
-A Poisson $\mbox{INGARCH}(p,q)$ (with $p,q\in\mathbb{N}_{0}$) model $\{Y_{t}\}_{t\in\mathbb{N}}$ is defined by
+The INGARCH nomenclature emerges from the fact that for a Poisson distribution (this assumption is imposed in the aforementioned papers), the mean equals its variance.
+Therefore modeling of mean implies modeling of variance like in the classic GARCH models introduced by Bollerslev (1986).
+Consequently, such models are considered in the literature as an integer counterpart of the GARCH models although this terminology should be used cautiously because ordinary GARCH processes are employed for modeling time-varying variances.
+A Poisson $\mbox{INGARCH}(p,q)$ (with $p,q\in\mathbb{N}$) model $\{Y_{t}\}_{t\in\mathbb{Z}}$ is defined by
 $Y_{t}|\mathcal{F}_{t-1}\sim\mbox{Poisson}(\lambda_{t}),\quad\lambda_{t}\equiv E(Y_{t}|\mathcal{F}_{t-1})=\beta_{0}+\sum_{i=1}^{p}\beta_{i}Y_{t-i}+\sum_{j=1}^{q}\alpha_{j}\lambda_{t-j},$ (1)
-where $\mathcal{F}_{t-1}=\sigma\{Y_{t-1},\ldots,Y_{0},\lambda_{0}\}$, $\beta_{0}>0$, $\beta_{i}\geq 0$ and $\alpha_{j}\geq 0$, for all $i=1,\ldots,p$ and $j=1,\ldots,q$.
+where $\mathcal{F}_{t}=\sigma\{Y_{s},s\leq t\}$, $\beta_{0}>0$, $\beta_{i}\geq 0$ and $\alpha_{j}\geq 0$, for all $i=1,\ldots,p$ and $j=1,\ldots,q$.
 Since $E(Y_{t}|\mathcal{F}_{t-1})=\mbox{Var}(Y_{t}|\mathcal{F}_{t-1})=\lambda_{t}$, the model specifies dynamic processes for both the conditional mean and variance, as previously stated.
@@
 However, due to limitations for fitting adequately real count time series data (e.g.
-not capturing all the sources of overdispersion although the model itself is overdispersed, inflation or deflation of zeros), other variants have been proposed based on such as the negative binomial INGARCH process by Zhu (2011) and Christou and Fokianos (2014) or the more general mixed Poisson INGARCH models by Christou and Fokianos (2015), Silva and Barreto-Souza (2019), among others.
-Zero-inflated versions of the Poisson and negative binomial INGARCH models were proposed by Zhu (2012a), while processes dealing with both overdispersion and underdispersion were proposed by Zhu (2012b, c) and Xu et al.
+not capturing all the sources of overdispersion although the model itself is overdispersed), several variants of the Poisson INGARCH model have been proposed.
+These include the negative binomial (NB) INGARCH process by Zhu (2011) and Christou and Fokianos (2014) or the more general mixed Poisson INGARCH models by Christou and Fokianos (2015), Silva and Barreto-Souza (2019), among others.
+Zero-inflated versions of the Poisson and negative binomial INGARCH models have been studied by Zhu (2012a), while processes dealing with both overdispersion and underdispersion were proposed by Zhu (2012b, c) and Xu et al.
 (2012).
-An alternative process to the linear model for the conditional mean, as in Equation (1), was introduced by Fokianos and Tjøstheim (2011) through the log-linear INGARCH processes.
-These models cope with both negative and positive autocorrelation function and allow inclusion of covariates in a straightforward way.
-Another different approach was recently proposed by Weiß et al.
-(2022), where a softplus link function is assumed instead of a logarithmic one.
-Based on the discussion so far, the Poisson and Negative Binomial distributions are frequently used in applications.
-In several cases, the Negative Binomial model fits more adequate due to the fact that this distribution depends on a dispersion parameter which facilitates flexible modeling.
+A negative binomial quasi-likelihood based inference for general integer-valued time series models with application to the Poisson and NB INGARCH models has been developed by Aknouche et al.
+(2018).
+An alternative process to the linear model given in (1), was introduced by Fokianos and Tjøstheim (2011) through the log-linear INGARCH processes.
+These models cope with both negative and positive autocorrelation functions and allow the inclusion of covariates in a straightforward way.
+More recently Weiß et al.
+(2022) considered a softplus link function instead of a logarithmic one.
+Based on the discussion so far, the Poisson and negative binomial distributions are frequently used in applications for count time series analysis.
+In several cases, the negative binomial model fits more adequately because this distribution depends on a dispersion parameter which facilitates flexible modeling.
 The dispersion parameter is usually assumed fixed and is estimated by using the Pearson residuals; see Christou and Fokianos (2014) for more.
-The main goal of this paper is to propose a novel and general class of INGARCH processes based on the mixed Poisson distributions (the Negative Binomial distribution falls in this class) by allowing time-varying mean and dispersion parameters.
+The main goal of this paper is to propose a novel and general class of INGARCH processes based on the mixed Poisson distributions (the negative binomial distribution belongs to this class) by allowing both time-varying mean and dispersion parameters.
 We call these models time-varying dispersion INGARCH (tv-DINGARCH) models.
-Hence, we develop methodology and study new models for integer-valued time series, where the assumption of constant dispersion might not be hold.
-The advantage of the tv-DINGARCH processes over the ordinary INGARCH models will be illustrated in Section 5, where we show that such processes fit better real data whose stylistic facts cannot be dealt with the existing INGARCH models.
+Hence, we develop methodology and introduce new models for integer-valued time series, where the assumption of constant dispersion might not hold.
+The usefulness of the tv-DINGARCH processes over the ordinary INGARCH models will be illustrated in Section 5, where we show that such processes fit better real data whose stylistic facts cannot be dealt with the existing INGARCH models.
 For example, in a count regression context, Barreto-Souza and Simas (2016) demonstrated through a data application on the attendance behavior of high school juniors that a constant dispersion assumption can be violated.
 Generalized linear models allowing regression structures for both mean and dispersion/precision parameters have been discussed previously by Efron (1986) and Smyth (1989), among others.
-In this work, we develop further this line of research by studying in detail joint modeling of mean and dispersion parameters in the context of dependent data.
-We put special emphasis in the case of INGARCH models for count time series and demonstrate the usefulness of such methodological development.
-Besides exploring the traditional INGARCH models, the proposed class contains the case that admits constant mean and time-dependent variance, similar to the spirit of the ordinary GARCH models.
-Another important feature is volatility modeling, which is a well-explored topic for the case of continuous valued time series but has been neglected in the context of integer-valued time series.
-Although the traditional Poisson INGARCH processes consider a time-dependent conditional variance, this in turn is driven by the dynamics of the mean process.
-The main disadvantage of this approach is that imposes severe restrictions as illustrated in the real data application in this paper.
-The tv-DINGARCH models relax such an assumption by considering a time-dependent dispersion parameter, therefore also controlling the conditional variance and allowing for additional source of volatility.
+In this work, we further develop this line of research by studying in detail the joint modeling of mean and dispersion parameters in the context of dependent data.
+We put special emphasis on the case of INGARCH models for count time series and demonstrate the usefulness of such methodological development.
+Besides extending the traditional INGARCH models, the proposed class includes the case of constant mean and time-dependent variance, which is a new model class that does not fall in the ordinary INGARCH model approach.
+Another important feature is volatility modeling, which is a well-explored topic for the case of continuous-valued time series but has been neglected in the context of integer-valued time series.
+Although the traditional Poisson INGARCH processes consider a time-dependent conditional variance, this in turn is solely driven by the dynamics of the mean process.
+The main disadvantage of this approach is that imposes severe restrictions on the volatility modelling as illustrated in the real data application in this paper.
+The class tv-DINGARCH models relaxes this assumption by considering a time-dependent dispersion parameter, therefore also controlling the conditional variance and allowing for additional sources of volatility.
 In the context of time-varying models for INGARCH processes, Roy and Karmakar (2021) introduced a Poisson INGARCH model with time-varying coefficients and Ratnayake and Samaranayake (2023) study a time-varying INGARCH model based on the zero-inflated Poisson distribution.
@@
 In addition, we consider dynamics that enter through the dispersion parameter.
-One advantage of the latter feature is that we are able to establish desirable stability properties for the models we consider, such as stationarity and ergodicity, based on e-chain theory.
+One advantage of the latter feature is that we are able to establish desirable stability properties for the models we consider, such as stationarity and ergodicity, by employing e-chain theory; see Meyn and Tweedie (1993).
 The paper is organized as follows.
@@
 Some technical results and proofs are provided in the Appendix.
-This paper contains Supplementary Material, which is available upon request.
+This paper contains Supplementary Material.
 Notation:
 We say that a random variable $Y$ follows a mixed Poisson (MP) distribution if satisfies the stochastic representation $Y|Z=z\sim\mbox{Poisson}(\lambda z)$, with $Z$ following some non-negative distribution with $E(Z)=1$ (standardization) and $\mbox{Var}(Z)=\phi^{-1}$, for $\lambda,\phi>0$.
-We denote $Y\sim\mbox{MP}(\lambda,\phi)$.
 In this case, $E(Y)=\lambda$ and $\mbox{Var}(Y)=\lambda+\lambda^{2}/\phi$.
+We denote $Y\sim\mbox{MP}(\lambda,\phi;\mathcal{D})$, where $\mathcal{D}$ denotes the distribution of $Z$.
 A random variable $Z$ following a Gamma distribution, with respective shape and scale parameters $a>0$ and $b>0$, is denoted by $Z\sim\mbox{Gamma}(a,b)$, where $E(Z)=a/b$ and $\mbox{Var}(Z)=a/b^{2}$.
-If the mixing variable $Z\sim\mbox{Gamma}(\phi,\phi)$, then we obtain that $Y$ follows a negative binomial (NB) distribution, i.e.
+If the mixing variable $Z\sim\mbox{Gamma}(\phi,\phi)$ (i.e.
+$\mathcal{D}=\mbox{Gamma}$), then we obtain that $Y$ follows a negative binomial (NB) distribution, i.e.
 $Y\sim\mbox{NB}(\lambda,\phi)$.
@@
 The induced $p$-norm for a $m\times l$ matrix ${\bf C}$ is then defined by $\|{\bf C}\|_{p}=\max_{{\bf x}\neq{\bf 0}}\{\|{\bf Cx}\|_{p}/\|{\bf x}\|_{p}:\,{\bf x}\in\mathbb{R}^{l}\}$, for $p\in[1,\infty]$.
+For a square matrix A, we denote its spectral radius by $\rho({\bf A})$.
 2 The tv-DINGARCH Models
-In this section we propose a class of time-varying dispersion INGARCH models by allowing both the mean and the dispersion parameter to depend on lagged valued of the observed series and their past values as follows.
+In this section, we propose a class of time-varying dispersion INGARCH models by allowing both mean and dispersion parameters to depend on the lagged values of the observed series and their past values as follows.
 Definition 2.1.
@@
 $\lambda_{t}=f(Y_{t-1},\ldots,Y_{t-p_{1}};\lambda_{t-1},\ldots,\lambda_{t-q_{1}}),\quad\phi_{t}=g(Y_{t-1},\ldots,Y_{t-p_{2}};\phi_{t-1},\ldots,\phi_{t-q_{2}}),$ (2)
-where $\mathcal{F}_{t-1}=\sigma\{Y_{t-1},\ldots,Y_{0},\phi_{0},\lambda_{0}\}$, and $(\lambda_{0},\phi_{0})$ denoting some starting value, $f:\mathbb{N}^{p_{1}}\times(0,\infty)^{q_{1}}\rightarrow(0,\infty)$, and $g:\mathbb{N}^{p_{2}}\times(0,\infty)^{q_{2}}\rightarrow(0,\infty)$.
+where $\mathcal{F}_{t}=\sigma\{Y_{s},s\leq t\}$ and $f:\mathbb{N}^{p_{1}}\times(0,\infty)^{q_{1}}\rightarrow(0,\infty)$, and $g:\mathbb{N}^{p_{2}}\times(0,\infty)^{q_{2}}\rightarrow(0,\infty)$.
 Of particular interest to our study and to diverse applications is the negative binomial (NB) model when it is imposed in Definition 2.1.
@@
 $P(Y_{t}=y|\mathcal{F}_{t-1})=\dfrac{\Gamma(y+\phi_{t})}{y!\Gamma(\phi_{t})}\left(\dfrac{\lambda_{t}}{\lambda_{t}+\phi_{t}}\right)^{y}\left(\dfrac{\phi_{t}}{\lambda_{t}+\phi_{t}}\right)^{\phi_{t}},\quad y\in\mathbb{N}_{0}.$
-The negative binomial model will be assumed to prove finiteness of moments and to derive the asymptotic theory for conditional maximum likelihood estimators in what follows.
-Some of results hold in general for mixed Poisson models, see Theorem 2.1, which states stationarity and ergodicity of time-varying dispersion INGARCH models.
-Definition 2.1 is general and some additional assumptions on the function $f$ and $g$ are necessary to study in detail these processes.
-We consider the first order model ($p_{1}=p_{2}=q_{1}=q_{2}=1$) linear tv-DINGARCH processes, which will be defined in what follows.
-This particular linear parametric form is a common choice considered in the literature but also enables a thorough study of stability properties of processes defined by Definition 2.1.
+The negative binomial model will be assumed to prove the finiteness of moments and to derive the asymptotic theory for conditional maximum likelihood estimators in what follows.
+Definition 2.1 is general and some additional assumptions on the functions $f$ and $g$ are necessary to study in detail these processes.
+We consider the first-order model ($p_{1}=p_{2}=q_{1}=q_{2}=1$) linear tv-DINGARCH processes, which will be defined in what follows.
+This particular linear parametric form is a common choice considered in the literature but also enables a thorough study of stability properties of processes given by Definition 2.1.
 Definition 2.2.
@@
 Remark 2.1.
-Note that we are using $Y_{t-1}$ to model $\phi_{t}$ in eq.
+We use $Y_{t-1}$ to model $\phi_{t}$ in eq.
 (3), and more generally in Def.
 2.1, instead of terms such as $Y^{2}_{t-1}$ and $|Y_{t-1}|$ considered for the continuous GARCH models and their modification.
-This is so because $Y_{t-1}$ is non-negative in our case (count time series).
+This is so because $Y_{t-1}$ is non-negative as we deal with count data.
 Remark 2.2.
-Some particular cases are obtained from the linear tv-DINGARCH class defined above.
-For $\alpha_{1}=\alpha_{2}=0$, we obtain the ordinary linear mixed Poisson INGARCH (Christou and Fokianos, 2015; Silva and Barreto-Souza, 2019) models as particular cases.
+Some particular cases are obtained from the tv-DINGARCH class defined above.
+When $\alpha_{1}=\alpha_{2}=0$, we obtain the ordinary linear mixed Poisson INGARCH (Christou and Fokianos, 2015; Silva and Barreto-Souza, 2019) models as particular cases.
 Additionally, by taking $\alpha_{0}\rightarrow 0^{+}$, we also obtain the Poisson INGARCH model as a limiting member of our proposed class.
@@
 Under this setting, the mean of the INGARCH process is constant and the variance is time-dependent as in the case of ordinary GARCH models (Bollerslev, 1986).
-Such property does not hold for the standard Poisson INGARCH model (1).
-In this case, we refer to (3) as the Pure INGARCH (P-INGARCH) process.
-We use the term “pure" to connect the fact that our model mimics the traditional continuous GARCH models (constant mean and time-varying variance).
+Such property does not hold for ordinary Poisson INGARCH models given by Equation (1).
+When $\beta_{1}=\beta_{2}=0$ we call (3) Pure INGARCH (P-INGARCH) process.
+We use the term “pure" to emphasize the fact that the model mimics the traditional continuous GARCH models (constant mean and time-varying variance).
 Hence, our approach is general and encompasses many different models studied earlier in the literature.
-Some simulated trajectories of the linear tv-DINGARCH models for some parameter settings are shown in the Supplementary Material.
+Simulated trajectories of the linear tv-DINGARCH models for a few parameter settings are shown in the Supplementary Material.
 2.1 Stationarity and Ergodicity
 […]
-Conditions for the existence and stationarity of the process (3) can be established, for example, by following the strategy by Christou and Fokianos (2014), which relies on establishing weak dependence by Doukhan and Wintenberger (2008).
-We here obtain such desirable properties based on another approach, the e-chains theory (Meyn and Tweedie, 1993), as follows.
-The dynamic latent processes $\{\lambda_{t}\}_{t\geq 1}$ and $\{\phi_{t}\}_{t\geq 1}$ given in (3) can be rewritten in a matrix form.
-By defining $\boldsymbol{\xi}_{t}=(\lambda_{t},\phi_{t})^{\top}$, for $t\geq 1$, $\boldsymbol{\tau}=(\beta_{0},\alpha_{0})^{\top}$ ${\bf A}=\begin{pmatrix}\beta_{2}&0\\ 0&\alpha_{2}\end{pmatrix}$, and ${\bf B}=\begin{pmatrix}\beta_{1}&0\\ 0&\alpha_{1}\end{pmatrix}$, we have that $\boldsymbol{\xi}_{t}=\boldsymbol{\tau}+{\bf B}(Y_{t-1},Y_{t-1})^{\top}+{\bf A}\boldsymbol{\xi}_{t-1}$.
-Under this matrix representation, we can adapt the strategy in Liu (2012) to establish the existence and stationarity of our process.
+The dynamic latent processes $\{\lambda_{t}\}_{t\in\mathbb{Z}}$ and $\{\phi_{t}\}_{t\in\mathbb{Z}}$ given in (3) can be written compactly in a vector form.
+By defining $\boldsymbol{\xi}_{t}=(\lambda_{t},\phi_{t})^{\top}$, $\boldsymbol{\tau}=(\beta_{0},\alpha_{0})^{\top}$, ${\bf A}=\begin{pmatrix}\beta_{2}&0\\ 0&\alpha_{2}\end{pmatrix}$ and ${\bf B}=\begin{pmatrix}\beta_{1}&0\\ 0&\alpha_{1}\end{pmatrix}$, we obtain that $\boldsymbol{\xi}_{t}=\boldsymbol{\tau}+{\bf B}(Y_{t-1},Y_{t-1})^{\top}+{\bf A}\boldsymbol{\xi}_{t-1}$.
+Using this representation, we adapt the strategy in Liu (2012) to establish existence and stationarity of (3).
 In that work, the author provided stochastic properties of a bivariate Poisson INGARCH model.
-The key point is to show that $\{\boldsymbol{\xi}_{t}\}_{t\in\mathbb{N}}$ is an e-chain Meyn and Tweedie (1993, Ch.18).
-The following theorem holds true for any mixed Poisson model including the Negative Binomial model.
+The key point is to show that $\{\boldsymbol{\xi}_{t}\}_{t\in\mathbb{Z}}$ is an e-chain; see Meyn and Tweedie (1993, Ch.18).
+In addition, the property of Geometric Moment Contraction (GMC), as studied by Diaconis and Freedman (1999) and Wu and Shao (2004), enables us to prove the desired ergodicity and uniqueness of the stationary distribution of $\{\boldsymbol{\xi}_{t}\}_{t\in\mathbb{Z}}$.
+Remark 2.3.
+In what follows, we prove all theoretical results for the negative binomial linear tv-DINGARCH models as given in 2.2 except Thm.
+2.1 and Corollary 2.1 which hold true for any linear tv-DINGARCH mixed Poisson model including the negative binomial model.
 Theorem 2.1.
-Let $\{Y_{t}\}_{t\in\mathbb{N}}$ be a tv-DINGARCH process as in (3).
-If there exists $p\in[1,\infty]$ such that $\|{\bf A}\|_{p}+2^{1-1/p}\|{\bf B}\|_{p}<1$, then the trivariate process $\{(Y_{t},\lambda_{t},\phi_{t})\}_{t\in\mathbb{N}}$ has a unique stationary and ergodic solution.
+Let $\{Y_{t}\}_{t\in\mathbb{Z}}$ be a tv-DINGARCH process (with fixed mixing distribution $\mathcal{D}$) as in (3).
+(a) If $\rho({\bf A}+{\bf B})<1$, then there exists at least one stationary distribution to $\{\boldsymbol{\xi}_{t}\}_{t\in\mathbb{Z}}$, where $\rho(\cdot)$ denotes the spectral radius.
+In addition, if $\|{\bf A}\|_{p}<1$ for some $1\leq p\leq\infty$, then the distribution is unique.
+(b) If $\|{\bf A}\|_{p}+2^{1-1/p}\|{\bf B}\|_{p}<1$ for some $p\in[1,\infty]$, then $\{\boldsymbol{\xi}_{t}\}_{t\in\mathbb{Z}}$ is a GMC Markov chain with a unique stationary and ergodic distribution.
 […]
 $\lambda_{t}=\beta_{0}+\sum_{i=1}^{p_{1}}\beta_{1i}Y_{t-i}+\sum_{j=1}^{p_{2}}\beta_{2j}\lambda_{t-j},\quad\phi_{t}=\alpha_{0}+\sum_{i=1}^{q_{1}}\alpha_{1i}Y_{t-i}+\sum_{j=1}^{q_{2}}\alpha_{2j}\phi_{t-j}.$ (4)
-Then by considering $\boldsymbol{\xi}_{t}=(\lambda_{t},\phi_{t})^{\top}=\boldsymbol{\tau}+\sum_{j=1}^{q}{\bf B}_{j}(Y_{t-j},Y_{t-j})^{\top}+\sum_{i=1}^{m}{\bf A}_{i}\boldsymbol{\xi}_{t-i}$, where $\{{\bf A}_{i}\}_{i=1}^{m}$ and $\{{\bf B}_{j}\}_{j=1}^{q}$ are
+If we denote by $\boldsymbol{\xi}_{t}=(\lambda_{t},\phi_{t})^{\top}=\boldsymbol{\tau}+\sum_{j=1}^{q}{\bf B}_{j}(Y_{t-j},Y_{t-j})^{\top}+\sum_{i=1}^{m}{\bf A}_{i}\boldsymbol{\xi}_{t-i}$, where $\{{\bf A}_{i}\}_{i=1}^{m}$ and $\{{\bf B}_{j}\}_{j=1}^{q}$ are defined by
 ${\bf A}_{j}=\begin{pmatrix}\beta_{2j}&0\\ 0&\alpha_{2j}\end{pmatrix}\leavevmode\nobreak\ \mbox{and}\leavevmode\nobreak\ {\bf B}_{j}=\begin{pmatrix}\beta_{1j}&0\\ 0&\alpha_{1j}\end{pmatrix},$
-$m=\max(p_{1},p_{2})$, and $q=\max(q_{1},q_{2})$. Then, following the same steps as in the proof of Theorem 2.1 the following result can be established.
+$m=\max(p_{1},p_{2})$, and $q=\max(q_{1},q_{2})$, then, following the same steps as in the proof of Theorem 2.1, the next result can be established; see Liu (2012, Prop. 4.3.1)
 Corollary 2.1.
 […]
-We close this section by proving finiteness of the moments (see Appendix for a proof) of the first-order negative binomial INGARCH process, which will be used in the next section to establish the asymptotic normality of the conditional maximum likelihood estimators.
+We close this section by proving the finiteness of the moments (see Appendix for a proof) of the first-order negative binomial tv-DINGARCH process.
 Theorem 2.2.
-For $\|{\bf A}\|_{1}+\|{\bf B}\|_{1}<1$, the first-order negative binomial INGARCH process $\{Y_{t}\}_{t\in\mathbb{N}}$ given by (3) has finite moments of order $k\in\mathbb{N}$, for an arbitrary $k$.
+If $\|{\bf A}\|_{1}+\|{\bf B}\|_{1}<1$, then the first-order negative binomial tv-DINGARCH process $\{Y_{t}\}_{t\in\mathbb{Z}}$ given by (3) satisfies that $E|Y_{t}|^{k}<\infty$, $E|\lambda_{t}|^{k}<\infty$ and $E|\phi_{t}|^{k}<\infty$ for any $k\in\mathbb{N}$.
+Remark 2.4.
+We conjecture that Theorem 2.2 can be extended to hold for any linear mixed Poisson tv-DINGARCH model by combining our arguments and the results of Karlis and Xekalaki (2005, Subsection 3.2) concerning moments of mixed Poisson distributions.
+This is out of the scope of this work since the focus is on the negative binomial model.
 3 Estimation and Asymptotic Theory
 […]
-3.1 Estimators and Asymptotics
-Let $\boldsymbol{\theta}=(\theta_{1},\ldots,\theta_{6})^{\top}=(\beta_{0},\beta_{1},\beta_{2},\alpha_{0},\alpha_{1},\alpha_{2})^{\top}$ be the parameter vector, and $y_{1},\ldots,y_{n}$ be a realization of a NB tv-DINGARCH process $\{Y_{t}\}_{t=1}^{n}$.
-The conditional log-likelihood function of $Y_{2},\ldots,Y_{n}$ given $Y_{1}=y_{1}$ is given by $\ell(\boldsymbol{\theta})=\sum_{t=2}^{n}\ell_{t}(\boldsymbol{\theta})$, where
-$\ell_{t}(\boldsymbol{\theta})=y_{t}[\log\lambda_{t}-\log(\lambda_{t}+\phi_{t})]+\phi_{t}[\log\phi_{t}-\log(\lambda_{t}+\phi_{t})]+\log\Gamma(y_{t}+\phi_{t})-\log\Gamma(\phi_{t})-\log y_{t}!,$
-for $t=2,\ldots,n$, where we have omitted the dependence of $\lambda_{t}$ and $\phi_{t}$ on $\boldsymbol{\theta}$ for simplicity of notation.
-In practice, it is necessary to set some initial values for $\lambda_{1}$ and $\phi_{1}$, which are fixed in this paper.
-More specifically, we get such initial values based on the two first empirical moments of the count time series.
-This strategy has worked well as demonstrated in the simulated results to be discussed in the sequence.
-The conditional maximum likelihood estimator (CMLE) of $\boldsymbol{\theta}$, say $\widehat{\boldsymbol{\theta}}$, is given by $\widehat{\boldsymbol{\theta}}=\mbox{argmax}_{\boldsymbol{\theta}\in\Theta}\,\ell(\boldsymbol{\theta})$, where $\Theta=(0,\infty)\times[0,\infty)^{2}\times(0,\infty)\times[0,\infty)^{2}$ denotes the parameter space.
-The score function associated to the conditional log-likelihood function is ${\bf U}(\boldsymbol{\theta})=\partial\ell(\boldsymbol{\theta})/\partial\boldsymbol{\theta}=\sum_{t=2}^{n}{\bf U}_{t}(\boldsymbol{\theta})$, with
+Remark 3.1.
+Some earlier work, in the context of negative-binomial quasi-likelihood inference for count time series models was reported by Aknouche et al.
+(2018).
+More precisely these authors studied model (3) but with a constant dispersion parameter.
+There are several differences between our work and that of Aknouche et al.
+(2018).
+  1. 1.
+    The first major difference is the type of models considered for the dispersion parameter.
+We assume a dynamic model–recall equation (3)–whereby the choice $\alpha_{1}=\alpha_{2}=0$ implies the model considered by Aknouche et al.
+(2018).
+So, in this sense, we extend this contribution.
+  2. 2.
+    The estimation method considered by Aknouche et al.
+(2018) is different from the direct maximization problem we are studying in this work.
+Indeed, Aknouche et al.
+(2018) consider a two-stage NB-QMLE (see their Sec.
+3.2).
+We have developed theory for direct maximization of the negative binomial quasi-likelihood under an extended model.
+  3. 3.
+    In Aknouche et al.
+(2018) the invertibility of the Fisher information matrix is not proved but instead it is assumed; see Assumption (A9) (existence, finiteness, and invertibility of $J_{r}$ following the notation of their paper).
+The invertibility of the information matrix is addressed by our contribution, in a more general form, by the proof of Theorem 3.1.
+  4. 4.
+    Furthermore, these authors assume finiteness of moments (Aknouche et al., 2018, Assumption A3), which requires the existence of $\delta>1$ such that $E(X_{t}^{\delta})<\infty$.
+These are natural conditions to be assumed since the authors are studying exclusively quasi-likelihood inference.
+Instead we have proved such condition for this extended model we consider.
+There are additional differences, see for instance (Aknouche et al., 2018, Assumption A2) and compare with our Lemma A-1.
+3.1 Likelihood Inference
+Denote by $\boldsymbol{\theta}_{1}=(\beta_{0},\beta_{1},\beta_{2})^{\top}$, $\boldsymbol{\theta}_{2}=(\alpha_{0},\alpha_{1},\alpha_{2})^{\top}$, and $\boldsymbol{\theta}=(\boldsymbol{\theta}_{1}^{\top},\boldsymbol{\theta}_{2}^{\top})^{\top}=(\theta_{1},\ldots,\theta_{6})^{\top}$ the parameter vector and suppose that $Y_{1},\ldots,Y_{n}$ is a sample from a NB tv-DINGARCH process $\{Y_{t}\}_{t=1}^{n}$.
+The conditional log-likelihood function of $Y_{2},\ldots,Y_{n}$ given $Y_{1}=y_{1}$ is given by (up to a constant)
+$\ell(\boldsymbol{\theta})$ $=\sum_{t=2}^{n}\ell_{t}(\boldsymbol{\theta})$ (5)
+$\ell_{t}(\boldsymbol{\theta})$ $=y_{t}[\log\lambda_{t}-\log(\lambda_{t}+\phi_{t})]+\phi_{t}[\log\phi_{t}-\log(\lambda_{t}+\phi_{t})]+\log\Gamma(y_{t}+\phi_{t})-\log\Gamma(\phi_{t}),$
+for $t=2,\ldots,n$, where we have omitted the dependence of $\lambda_{t}$ and $\phi_{t}$ on $\boldsymbol{\theta}$ for simplicity.
+The computation of the log-likelihood function (5) depends on the initial values of $\lambda_{0}$ and $\phi_{0}$.
+If we denote by $\tilde{\lambda}_{0}$ and $\tilde{\phi}_{0}$ some fixed or random starting values of the process (3), then the observed log-likelihood function is given by
+~$\tilde{\ell}(\boldsymbol{\theta})$ $=\sum_{t=1}^{n}\tilde{\ell}_{t}(\boldsymbol{\theta})$ (6)
+~$\tilde{\ell}_{t}(\boldsymbol{\theta})$ $=y_{t}[\log\tilde{\lambda}_{t}-\log(\tilde{\lambda}_{t}+\tilde{\phi_{t})}]+\tilde{\phi}_{t}[\log\tilde{\phi}_{t}-\log(\tilde{\lambda}_{t}+\tilde{\phi}_{t})]+\log\Gamma(y_{t}+\tilde{\phi}_{t})-\log\Gamma(\tilde{\phi}_{t}),$
+with $\tilde{\lambda}_{t}=\beta_{0}+\beta_{1}Y_{t-1}+\beta_{2}\tilde{\lambda}_{t-1}$ and $\tilde{\phi}_{t}=\alpha_{0}+\alpha_{1}Y_{t-1}+\alpha_{2}\tilde{\phi}_{t-1}$.
+The conditional maximum likelihood estimator (CMLE) of $\boldsymbol{\theta}$, say $\widehat{\boldsymbol{\theta}}$, is given by $\widehat{\boldsymbol{\theta}}=\mbox{argmax}_{\boldsymbol{\theta}\in\Theta}\tilde{\ell}(\boldsymbol{\theta})$, where $\Theta=(0,\infty)\times[0,\infty)^{2}\times(0,\infty)\times[0,\infty)^{2}$ denotes the parameter space.
+In Lemma A-1 in the Appendix, we show that the log-likelihood function (6) is approximated by (5) with $(Y_{t},\lambda_{t},\phi_{t})_{t\in\mathbb{Z}}$ the stationary and ergodic process obtained by Theorem 2.1; see also Francq and Zakoïan (2004), Straumann and Mikosh (2006) and Christou and Fokianos (2014), among others.
+Note that (6) is the observed likelihood function but we will be employing (5) in what follows (including the proof of Theorem 3.1 about the asymptotic normality of the CMLEs) for simplicity of notation and without loss of generality.
+In practice, we obtain initial values by employing the two first empirical moments of the data.
+This strategy has worked well; see Sec.
+3.2 for empirical evidence.
+The score function associated with the conditional log-likelihood function is
+${\bf U}(\boldsymbol{\theta})=\frac{\partial\ell(\boldsymbol{\theta})}{\partial\boldsymbol{\theta}}=\sum_{t=2}^{n}{\bf U}_{t}(\boldsymbol{\theta}),$ (7)
+with
 ${\bf U}_{t}(\boldsymbol{\theta})=\left(S_{1t}\dfrac{\partial\lambda_{t}}{\partial\beta_{0}},S_{1t}\dfrac{\partial\lambda_{t}}{\partial\beta_{1}},S_{1t}\dfrac{\partial\lambda_{t}}{\partial\beta_{2}},S_{2t}\dfrac{\partial\phi_{t}}{\partial\alpha_{0}},S_{2t}\dfrac{\partial\phi_{t}}{\partial\alpha_{1}},S_{2t}\dfrac{\partial\phi_{t}}{\partial\alpha_{2}}\right)^{\top},$
@@
 Explicit expressions for the derivatives involving $\lambda_{t}$ and $\phi_{t}$ are presented in the Supplementary Material.
-To establish the asymptotic normality of the CMLEs, the following proposition is necessary.
+To establish the asymptotic normality of the CMLEs, the following lemma (its proof given in the Appendix) and proposition are necessary.
+Lemma 3.1.
+For $\|{\bf A}\|_{1}+\|{\bf B}\|_{1}<1$, the $t$-th term of the score function ${\bf U}_{t}(\boldsymbol{\theta})$ is integrable for all $t\geq 2$.
 Proposition 3.1.
@@
 Proof.
+We will first show that $E({\bf U}_{t}(\boldsymbol{\theta})|\mathcal{F}_{t-1})=0$.
+This fact and the integrability of ${\bf U}_{t}(\boldsymbol{\theta})$ stated in Lemma 3.1 will give us the desirable result.
 Since $E(Y_{t}|\mathcal{F}_{t-1})=\lambda_{t}$ by definition, we immediately obtain that $E(S_{1t}(\boldsymbol{\theta})|\mathcal{F}_{t-1})=0$.
-We now compute the conditional expectation of $S_{2t}$ given $\mathcal{F}_{t-1}$, which involves $E\left(\Psi(Y_{t}+\phi_{t})|\mathcal{F}_{t-1}\right)$.
+We compute the conditional expectation of $S_{2t}$ given $\mathcal{F}_{t-1}$, which involves the term $E\left(\Psi(Y_{t}+\phi_{t})|\mathcal{F}_{t-1}\right)$ and it is the most difficult to deal with.
 For $a>0$ and $|c|<1$, we have that
@@
 where $\Gamma^{\prime}(a)=d\Gamma(a)/da$. Using the above result, it follows that
-$E\left(\Psi(Y_{t}+\phi_{t})|\mathcal{F}_{t-1}\right)=\dfrac{(\lambda_{t}\phi_{t}^{-1}+1)^{-\phi_{t}}}{\Gamma(\phi_{t})}\sum_{y=0}^{\infty}\dfrac{\Gamma^{\prime}(y+\phi_{t})}{y!}\left(\dfrac{\lambda_{t}}{\lambda_{t}+\phi_{t}}\right)^{y}=\Psi(\phi_{t})-\log\left(\dfrac{\phi_{t}}{\lambda_{t}+\phi_{t}}\right),$
-and then we obtain that $E(S_{2t}(\boldsymbol{\theta})|\mathcal{F}_{t-1})=0$. Hence, for all $t\geq 1$, $E({\bf U}_{t}(\boldsymbol{\theta})|\mathcal{F}_{t-1})=0$.
-We will now show that ${\bf U}_{t}(\boldsymbol{\theta})$ is integrable.
-We have that $|S_{1t}|\leq|y_{t}/\lambda_{t}-1|\leq y_{t}/\beta_{0}+1$.
-Furthermore, $\partial\lambda_{t}/\partial\beta_{i}$ is a linear combination of $y_{t-1},\ldots,y_{1}$ and $\lambda_{1}$, for $i=0,1,2$.
-Using these results and under condition $\|{\bf A}\|_{1}+\|{\bf B}\|_{1}<1$, Theorem 2.2 gives that the score function associated with $\beta_{0},\beta_{1}$ and $\beta_{2}$ is integrable.
-We now study the score function associated with $\alpha_{0}$, $\alpha_{1}$ and $\alpha_{2}$.
-Note that $\left|-\dfrac{y_{t}-\lambda_{t}}{\lambda_{t}+\phi_{t}}\right|\leq\dfrac{y_{t}}{\lambda_{t}+\phi_{t}}+1\leq\dfrac{y_{t}}{2}\left(\dfrac{1}{\beta_{0}}+\dfrac{1}{\alpha_{0}}\right)+1$, where we have used that fact that $\dfrac{1}{a+b}\leq\dfrac{1}{2}\left(\dfrac{1}{a}+\dfrac{1}{b}\right)$ for $a,b>0$.
-Also, by using that $\dfrac{1}{2x}<\log x-\Psi(x)<\dfrac{1}{x}$ for $x>0$ (Alzer, 1997), we obtain that $0<\log\phi_{t}-\Psi(\phi_{t})<\phi_{t}^{-1}<\alpha_{0}^{-1}$.
-Similarly, it follows that $|\Psi(\lambda_{t}+\phi_{t})-\log(\lambda_{t}+\phi_{t})|<(\lambda_{t}+\phi_{t})^{-1}\leq\dfrac{1}{2}\left(\dfrac{1}{\beta_{0}}+\dfrac{1}{\alpha_{0}}\right)$.
-Therefore, $|S_{2t}|$ is bounded above by a linear combination of $y_{t}$.
-Moreover, $\dfrac{\partial\phi_{t}}{\partial\alpha_{j}}$ is a linear combination of $y_{t-1},\ldots,y_{1}$ and $\phi_{1}$, for $j=0,1,2$.
-Again, an application of Theorem 2.2 gives that the score function associated with $\alpha_{0}$, $\alpha_{1}$ and $\alpha_{2}$ is also integrable.
+$E\left(\Psi(Y_{t}+\phi_{t})|\mathcal{F}_{t-1}\right)=\dfrac{(\lambda_{t}\phi_{t}^{-1}+1)^{-\phi_{t}}}{\Gamma(\phi_{t})}\sum_{y=0}^{\infty}\dfrac{\Gamma^{\prime}(y+\phi_{t})}{y!}\left(\dfrac{\lambda_{t}}{\lambda_{t}+\phi_{t}}\right)^{y}=\Psi(\phi_{t})-\log\left(\dfrac{\phi_{t}}{\lambda_{t}+\phi_{t}}\right).$
+Thus
+$E(S_{2t}(\boldsymbol{\theta})|\mathcal{F}_{t-1})$ $=$ $E\left(-\dfrac{Y_{t}-\lambda_{t}}{\lambda_{t}+\phi_{t}}+\log\left(\dfrac{\phi_{t}}{\lambda_{t}+\phi_{t}}\right)+\Psi(y_{t}+\phi_{t})-\Psi(\phi_{t})\Big{|}\mathcal{F}_{t-1}\right)$
+$=$ $-\dfrac{E(Y_{t}|\mathcal{F}_{t-1})-\lambda_{t}}{\lambda_{t}+\phi_{t}}+\log\left(\dfrac{\phi_{t}}{\lambda_{t}+\phi_{t}}\right)+E(\Psi(y_{t}+\phi_{t})|\mathcal{F}_{t-1})-\Psi(\phi_{t})=0.$
+Hence, for all $t\geq 2$, $E({\bf U}_{t}(\boldsymbol{\theta})|\mathcal{F}_{t-1})=0$.
+From Lemma 3.1 and under the condition $\|{\bf A}\|_{1}+\|{\bf B}\|_{1}<1$, ${\bf U}_{t}(\boldsymbol{\theta})$ is integrable.
+Therefore the conclusion holds true.
 ∎
-Let us now discuss the asymptotic distribution of the CMLE for the NB tv-DINGARCH model.
+We next sketch the proof for obtaining the asymptotic distribution of the CMLE for the NB tv-DINGARCH model.
 A more detailed technical result will be established in Theorem 3.1.
-By using Proposition 3.1 and the square integrability of ${\bf U}(\boldsymbol{\theta})$ (this can be demonstrated using very similar arguments to the proof of Proposition 3.1), we can apply the Central Limit Theorem for martingale difference (for instance, see Corollary 3.1 from Hall and Heyde (1980)) to obtain that
+Recall (7).
+By using Proposition 3.1 and the square integrability of ${\bf U}(\boldsymbol{\theta})$ (this is proved using similar arguments to the proof of Proposition 3.1), we can apply the Central Limit Theorem for martingale difference (see Corollary 3.1 from Hall and Heyde (1980), for example) to obtain that
 $n^{-1/2}{\bf U}(\boldsymbol{\theta})\stackrel{{\scriptstyle d}}{{\longrightarrow}}N({\bf 0},\boldsymbol{\Omega}_{1}(\boldsymbol{\theta})),\quad\boldsymbol{\Omega}_{1}(\boldsymbol{\theta})\equiv\mbox{plim}_{n\rightarrow\infty}{\bf J}_{1}(\boldsymbol{\theta}),$ (8)
-as $n\rightarrow\infty$, with
-${\bf J}_{1}(\boldsymbol{\theta})\equiv\dfrac{1}{n}\sum_{t=2}^{n}\mbox{Var}({\bf U}_{t}(\boldsymbol{\theta})|\mathcal{F}_{t-1})=\dfrac{1}{n}\sum_{t=2}^{n}\begin{pmatrix}b_{t}\dfrac{\partial\lambda_{t}}{\partial\boldsymbol{\theta}_{*}}\dfrac{\partial\lambda_{t}}{\partial\boldsymbol{\theta}_{*}}^{\top}&{\bf 0}\\ {\bf 0}&l_{t}\dfrac{\partial\phi_{t}}{\partial\boldsymbol{\theta}^{*}}\dfrac{\partial\phi_{t}}{\partial\boldsymbol{\theta}^{*}}^{\top}\end{pmatrix},$
-$\boldsymbol{\theta}_{*}=(\theta_{1},\theta_{2},\theta_{3})^{\top}$, $\boldsymbol{\theta}^{*}=(\theta_{4},\theta_{5},\theta_{6})^{\top}$, $\Psi^{\prime}(x)=d\Psi(x)/dx$,
+as $n\rightarrow\infty$,
+${\bf J}_{1}(\boldsymbol{\theta})\equiv\dfrac{1}{n}\sum_{t=2}^{n}\mbox{Var}({\bf U}_{t}(\boldsymbol{\theta})|\mathcal{F}_{t-1})=\dfrac{1}{n}\sum_{t=2}^{n}\begin{pmatrix}b_{t}\dfrac{\partial\lambda_{t}}{\partial{\boldsymbol{\theta}}_{1}}\dfrac{\partial\lambda_{t}}{\partial{\boldsymbol{\theta}}_{1}}^{\top}&{\bf 0}\\ {\bf 0}&l_{t}\dfrac{\partial\phi_{t}}{\partial{\boldsymbol{\theta}}_{2}}\dfrac{\partial\phi_{t}}{\partial{\boldsymbol{\theta}}_{2}}^{\top}\end{pmatrix},$
+with
 $b_{t}=\dfrac{\phi_{t}}{\lambda_{t}(\lambda_{t}+\phi_{t})},\quad l_{t}=E\left(\Psi^{2}(Y_{t}+\phi_{t})|\mathcal{F}_{t-1}\right)-\dfrac{\lambda_{t}}{\phi_{t}(\lambda_{t}+\phi_{t})}-\left(\psi(\phi_{t})-\log\left(\dfrac{\phi_{t}}{\lambda_{t}+\phi_{t}}\right)\right)^{2},$
 for $t=2,\ldots,n$, where we have used that $E\left(Y_{t}\Psi(Y_{t}+\phi_{t})|\mathcal{F}_{t-1}\right)=\lambda_{t}\left\{\Psi(\phi_{t}+1)-\log\left(\dfrac{\phi_{t}}{\lambda_{t}+\phi_{t}}\right)\right\}$ and that $\Psi(x+1)=\Psi(x)+x^{-1}$, for $x>0$.
-Then, we apply the Law of Large Numbers for stationary and ergodic sequences to obtain that
+Then, we apply the Law of Large Numbers for stationary and ergodic sequences (see Thm. 6.28 of Breiman (1992)) to obtain that
 $n^{-1}\dfrac{\partial{\bf U}(\boldsymbol{\theta})}{\partial\boldsymbol{\theta}}\stackrel{{\scriptstyle p}}{{\longrightarrow}}\boldsymbol{\Omega}_{2}(\boldsymbol{\theta}),\quad\boldsymbol{\Omega}_{2}(\boldsymbol{\theta})\equiv\mbox{plim}_{n\rightarrow\infty}{\bf J}_{2}(\boldsymbol{\theta}),$ (9)
 as $n\rightarrow\infty$, where
-${\bf J}_{2}(\boldsymbol{\theta})\equiv\dfrac{1}{n}\sum_{t=2}^{n}E(-\nabla{\bf U}_{t}(\boldsymbol{\theta})|\mathcal{F}_{t-1})=\dfrac{1}{n}\sum_{t=2}^{n}\begin{pmatrix}b_{t}\dfrac{\partial\lambda_{t}}{\partial\boldsymbol{\theta}_{*}}\dfrac{\partial\lambda_{t}}{\partial\boldsymbol{\theta}_{*}}^{\top}&\bf 0\\ \bf 0&d_{t}\dfrac{\partial\phi_{t}}{\partial\boldsymbol{\theta}^{*}}\dfrac{\partial\phi_{t}}{\partial\boldsymbol{\theta}^{*}}^{\top}\end{pmatrix},$
+${\bf J}_{2}(\boldsymbol{\theta})\equiv\dfrac{1}{n}\sum_{t=2}^{n}E(-\nabla{\bf U}_{t}(\boldsymbol{\theta})|\mathcal{F}_{t-1})=\dfrac{1}{n}\sum_{t=2}^{n}\begin{pmatrix}b_{t}\dfrac{\partial\lambda_{t}}{\partial{\boldsymbol{\theta}}_{1}}\dfrac{\partial\lambda_{t}}{\partial{\boldsymbol{\theta}}_{1}}^{\top}&{\bf 0}\\ {\bf 0}&d_{t}\dfrac{\partial\phi_{t}}{\partial{\boldsymbol{\theta}}_{2}}\dfrac{\partial\phi_{t}}{\partial{\boldsymbol{\theta}}_{2}}^{\top}\end{pmatrix},$
 with
@@
 We have that $\widehat{\boldsymbol{\Sigma}}={\bf J}_{1}^{-1}(\widehat{\boldsymbol{\theta}})$ is a consistent estimator for $\boldsymbol{\Sigma}$.
-Another two consistent estimators for $\boldsymbol{\Sigma}$ are ${\bf S}^{-1}_{1}(\widehat{\boldsymbol{\theta}})$ and ${\bf S}^{-1}_{2}(\widehat{\boldsymbol{\theta}})$, where ${\bf S}_{1}(\boldsymbol{\theta})=n^{-1}\sum_{t=2}^{n}{\bf U}_{t}(\boldsymbol{\theta}){\bf U}_{t}(\boldsymbol{\theta})^{\top}$ and ${\bf S}_{2}(\boldsymbol{\theta})=-n^{-1}\sum_{t=2}^{n}\nabla{\bf U}_{t}(\boldsymbol{\theta})$.
-There are required additional technical conditions to ensure the consistency and asymptotic normality of the CML estimators, which are provided in the next theorem, where its proof can be found in the Appendix.
-Before doing that, we have to consider lower and upper bounds for the possible values that the parameters can assume and then we define
-=$\mathcal{B}=\{\boldsymbol{\theta}:\beta_{i,low}<\beta_{i}<\beta_{i,up},\,\,\alpha_{j,low}<\alpha_{j}<\alpha_{j,up},\,\,i,j=0,1,2\}.$ (8)
+There are required additional technical conditions to ensure the consistency and asymptotic normality of the CML estimators, which are provided in the next theorem, whose proof can be found in the Appendix.
+We first consider lower and upper bounds for the possible values of the components of the parameter vector by introducing (recall that $\theta_{3}=\beta_{2}$ and $\theta_{6}=\alpha_{2}$)
+=$\mathcal{B}=\Bigl{\{}\boldsymbol{\theta}:0<\theta_{i,low}\leq\theta_{i}\leq\theta_{i,up},i,=1,\ldots,6,\leavevmode\nobreak\ \mbox{and}\leavevmode\nobreak\ \theta_{3,up}<1,\theta_{6,up}<1\Bigr{\}}.$ (11)
 Theorem 3.1.
@@
 Then, there exists an open set $\mathcal{A}\subset\mathcal{B}$ such that the conditional log-likelihood function $\ell(\cdot)$ has a global maximum point on $\mathcal{A}$, say $\widehat{\boldsymbol{\theta}}$, with probability tending to 1 as $n\rightarrow\infty$.
-Furthermore, $\widehat{\boldsymbol{\theta}}$ is consistent for $\boldsymbol{\theta}$ and satisfy the asymptotic normality (7).
+Furthermore, $\widehat{\boldsymbol{\theta}}$ is consistent for $\boldsymbol{\theta}$ and satisfies (10).
 3.2 Monte Carlo Simulation
@@
 Table 1 summarises the results obtained for settings I and II.
-Empirical means and standard deviation (SD) of the CMLEs are provided by sample size.
-TResults show adequate performance of the estimation method, with empirical Monte Carlo means close to the values and standard deviation decreasing with the increase in sample size, as expected.
+Empirical means and standard deviation (SD) of the CMLEs are reported by sample size.
+The results show adequate performance of the estimation method, with empirical Monte Carlo means close to the values and standard deviation decreasing with the increase in sample size, as expected.
 Setting I $\beta_{0}=15$ $\alpha_{0}=0.5$ $\beta_{1}=0.2$ $\beta_{2}=0.25$ $\alpha_{1}=0.1$ $\alpha_{2}=0.3$
@@
 Dashed and dotted curves indicate densities estimated from the experiments carried out with $n=500$ and $n=1000$, respectively.
-The density curves of the parameter estimates are mostly overlapping, but some improvement can be seen with the increase of the sample size.
+The density curves of the parameter estimates are mostly overlapping, but some improvement can be seen with the increase in the sample size.
 Additional density plots for the other parameter configurations can be found in the Supplementary Material.
-Figure 1: Non-parametric density plots of standardized parameter estimates due to Setting I under sample sizes $n=500$ and $n=1000$.
-The solid line corresponds to the standard Gaussian density function.
+Figure 1: Non-parametric density plots of standardized parameter estimates due to Setting I when $n=500$ and $n=1000$. The solid line corresponds to the standard Gaussian density function.
 4 Testing Constant Dispersion
 Testing constant value of the dispersion parameter, i.e.
-$\phi_{t}=\alpha_{0}$ for all $t$ under model (3), is equivalent to testing $H_{0}:\alpha_{1}=\alpha_{2}=0$ against $H_{1}:\alpha_{1}\neq 0\,\,\mbox{or}\,\,\alpha_{2}\neq 0$.
-Note that the null hypothesis belongs to the boundary of the parameter space, which is a non-standard testing problem; for instance, see Self and Liang (1987), Andrews (2001), and Crainiceanu and Ruppert (2004).
-We develop and compare heuristically two parametric bootstrap methods; the classical or unrestricted, and the restricted bootstrap recently developed by Cavaliere et al.
+$\phi_{t}=\alpha_{0}$ for all $t$ under model (3), is equivalent to testing $H_{0}:\alpha_{1}=\alpha_{2}=0$ against $H_{1}:\alpha_{1}>0\,\,\mbox{or}\,\,\alpha_{2}>0$.
+Note that the null hypothesis implies that the parameters $\alpha_{1}$ and $\alpha_{2}$ belong to the boundary of the parameter space.
+Hence, this is a non-standard testing problem; see, for instance, Self and Liang (1987), Andrews (2001), and Crainiceanu and Ruppert (2004).
+We consider the likelihood ratio test to develop and compare heuristically two parametric bootstrap methods; the classical or unrestricted, and the restricted bootstrap recently developed by Cavaliere et al.
 (2016).
 The first method considers the usual parametric bootstrap (Efron and Tibshirani, 1993) replications based on the unrestricted CMLEs, while the latter uses the CMLEs under $H_{0}$.
-Algorithm 1 describes the estimation of the test’s p-value with $B$ replications of the restricted or unrestricted bootstrap.
+Algorithm 1 specifies computation of the likelihood ratio test’s p-value with $B$ replications of the restricted or unrestricted bootstrap.
 Input: $\boldsymbol{Y}$ observed count time series data
@@
 for $b\leftarrow 1:B$ do
-      3A) $\boldsymbol{Y}_{b}\sim\mbox{tv NB-INGARCH}(\widehat{\boldsymbol{\theta}}_{\mathcal{H}_{0}})$;
+      3A) $\boldsymbol{Y}_{b}\sim\mbox{NB tv-DINGARCH}(\widehat{\boldsymbol{\theta}}_{\mathcal{H}_{0}})$;
        // if restricted bootstrap
-      3B) $\boldsymbol{Y}_{b}\sim\mbox{tv NB-INGARCH}(\widehat{\boldsymbol{\theta}})$;
+      3B) $\boldsymbol{Y}_{b}\sim\mbox{NB tv-DINGARCH}(\widehat{\boldsymbol{\theta}})$;
        // if unrestricted bootstrap
       4.
-Obtain $\widehat{\boldsymbol{\theta}}_{\mathcal{H}_{0}}^{b}$ and $\boldsymbol{\widehat{\theta}}^{b}$ fitting tv NB-INGARCH models to $\boldsymbol{Y}_{b}$ under the null and alternative hypothesis;
+Obtain $\widehat{\boldsymbol{\theta}}_{\mathcal{H}_{0}}^{b}$ and $\boldsymbol{\widehat{\theta}}^{b}$ fitting NB tv-DINGARCH models to $\boldsymbol{Y}_{b}$ under the null and alternative hypothesis;
       5. Let $LR^{b}=-2(\ell(\widehat{\boldsymbol{\theta}}^{b}_{\mathcal{H}_{0}})-\ell(\boldsymbol{\widehat{\theta}}^{b}))$, the replicated LR statistic;
@@
 6. If $p_{B}=\sum_{b=1}^{B}I\{LR^{b}>LR\}/B<\alpha$ reject $\mathcal{H}_{0}$;
-Algorithm 1 Bootstrap likelihood ratio test of constant ($\mathcal{H}_{0}:\alpha_{1}=\alpha_{2}=0$) versus time-varying dispersion ($\mathcal{H}_{1}:\alpha_{1}\neq 0\mbox{ or }\alpha_{2}\neq 0$) for a tv NB-INGARCH model.
+Algorithm 1 Bootstrap likelihood ratio test of constant ($\mathcal{H}_{0}:\alpha_{1}=\alpha_{2}=0$) versus time-varying dispersion ($\mathcal{H}_{1}:\alpha_{1}>0\mbox{ or }\alpha_{2}>0$) for a NB tv-DINGARCH model.
 Alternatives to step 3 yield restricted (3A) or unrestricted bootstrap (3B) estimators of the test’s p-value, $p_{B}$, where $B$ is the number of replications.
 We use a Monte Carlo simulation study to investigate how these methods achieve the desirable significance levels.
-Time series from the NB tv-DINGARCH process with 200 time observations are simulated under the null hypothesis using four different (varying) settings of the parameter vector $(\beta_{0},\beta_{1},\beta_{2},\alpha_{0})^{\top}$ as follows: (C1) $\beta_{0}=2,\alpha_{0}=1$, $\beta_{1}=0.4,\beta_{2}=0.3$, (C2) $\beta_{0}=2,\alpha_{0}=1$, $\beta_{1}=0,\beta_{2}=0$, (C3) $\beta_{0}=3,\alpha_{0}=0.5$, $\beta_{1}=0.3,\beta_{2}=0.4$, and (C4) $\beta_{0}=3,\alpha_{0}=0.5$, $\beta_{1}=0,\beta_{2}=0$.
+Time series from the NB tv-DINGARCH process with 200 observations are simulated under the null hypothesis using four different (varying) settings of the parameter vector $(\beta_{0},\beta_{1},\beta_{2},\alpha_{0})^{\top}$ as follows: (C1) $\beta_{0}=2,\alpha_{0}=1$, $\beta_{1}=0.4,\beta_{2}=0.3$, (C2) $\beta_{0}=2,\alpha_{0}=1$, $\beta_{1}=0,\beta_{2}=0$, (C3) $\beta_{0}=3,\alpha_{0}=0.5$, $\beta_{1}=0.3,\beta_{2}=0.4$, and (C4) $\beta_{0}=3,\alpha_{0}=0.5$, $\beta_{1}=0,\beta_{2}=0$.
 For each configuration, 500 Monte Carlo replications are used to calculate the empirical significance levels by employing the competing methodologies.
@@
 Configuration
-Significance
+Nominal
 level
@@
 0.10 0.094 0.058
-Table 2: Nominal significance levels produced by the restricted and unrestricted bootstrap hypothesis tests for $H_{0}:\alpha_{1}=\alpha_{2}=0$ for various parameter configurations.
-Notably, the restricted parametric bootstrap agrees with the set significance levels, something that occurs under the four parameter configurations investigated in this study, while the unrestricted parametric bootstrap does not provide satisfactory results by underestimating the significance levels.
-These results show the importance of considering a restricted bootstrap for testing constant dispersion in the tv-DINGARCH models.
+Table 2: Achieved significance levels of likelihood ratio test employing Algorithm 1 for various parameter configurations.
+Notably, the restricted parametric bootstrap achieves levels close to nominal significance levels, something that occurs under all four parameter configurations investigated in this study.
+The unrestricted parametric bootstrap does not provide satisfactory results and it underestimates the significance levels.
+Hence, in applications, it is important to employ a restricted bootstrap for testing constant dispersion in the tv-DINGARCH models.
+The results obtained from both restricted and unrestricted bootstrap were already expected.
+As discussed by Cavaliere et al.
+(2016), it is well-known that the unrestricted (classical) bootstrap does not apply when one parameter belongs to the boundary of the parameter space.
+This is indeed the case where the null hypothesis is that $H_{0}:\alpha_{1}=\alpha_{2}=0$, both parameters belong to the boundary space.
+The restricted bootstrap proposed by Cavaliere et al.
+(2016) overcomes this issue and provides valid results as theoretically proven in their paper and empirically illustrated here for DINGARCH models.
 5 Measles Data Analysis
 The linear negative binomial tv-DINGARCH model is now applied to modeling the weekly number of reported measles infections in North Rhine-Westphalia, Germany.
-The series is observed between January 2001 and May 2013 (646 observations), publicly available in the R package tscount.
+The series is observed between January 2001 and May 2013 (646 observations).
+These data are publicly available in the R package tscount.
 Figure 2 displays $\{Y_{t}\}_{t=1}^{646}$ on the left, and the series autocorrelation function on the right.
 Figure 2: On the left, weekly cases of measles reported in North Rhine-Westphalia, Germany, between January 2001 and May 2013. The autocorrelation function of the series is shown on the right.
-Fitting of the NB tv-DINGARCH model will be compared to the fit of ordinary NBINGARCH model with an identical mean time series structure.
+The fitting of the NB tv-DINGARCH model will be compared to the fit of the ordinary NB INGARCH model with an identical mean time series structure.
 The purpose of this exercise is to assess how a time-varying dispersion changes the model adequacy to this real data.
-To this end, we will evaluate goodness-of-fit and predictions for each model.
+To this end, we will evaluate goodness-of-fit and predictions for each model, i.e.
 $\mathcal{M}_{tv}$ will denote the NB tv-DINGARCH model while $\mathcal{M}_{ord}$ denotes the ordinary case (which assumes a constant dispersion).
@@
 $\alpha_{2}$ 0.000 (0.010) (0.000, 0.020)
-Table 3: Conditional maximum likelihood estimates, standard errors (se), and 95% confidence intervals for log-linear tv-DINGARCH and ordinary log-linear INGARCH models applied to weekly counts of measles in Germany.
-In addition, we consider Probability Integral Transform (PIT) plots (Czado et al., 2009), an approach that enables the comparison of count data models via their predictive distributions.
+Table 3: Conditional maximum likelihood estimates, standard errors (se), and 95% confidence intervals for $\mathcal{M}_{tv}$ and $\mathcal{M}_{ord}$ models applied to weekly counts of measles in Germany.
+In addition, we consider Probability Integral Transform (PIT) plots (Czado et al., 2009), an approach that enables the comparison of count time series models via their predictive distributions.
 A model providing a good fit to the data in this aspect will render a PIT plot resembling a uniform distribution, where major deviations typically indicate problems of overdispersion or underdispersion of the model’s predictive distribution.
-These are reported in Figure 3 for both models, and while that of $\mathcal{M}_{tv}$ is near the uniform as desired, the upside-down U shaped PIT of $\mathcal{M}_{ord}$ points to a non adequate fit.
-Figure 3: Probability Integral Transform (PIT) assessment for $\mathcal{M}_{ord}$ (on the left) and $\mathcal{M}_{tv}$ (to the right).
+These graphs are shown in Figure 3 for both models and while the plot for $\mathcal{M}_{tv}$ resembles uniformity, as desired, the upside-down U-shaped PIT of $\mathcal{M}_{ord}$ indicates a non adequate fit.
+Figure 3: PIT plots for $\mathcal{M}_{ord}$ (left) and $\mathcal{M}_{tv}$ (right).
 It is often of interest to practitioners to select between models according to their forecasting power.
-To that end, we consider a forecasting exercise that explroes one-step-ahead (OSA) prediction from $\mathcal{M}_{tv}$ and $\mathcal{M}_{ord}$.
-This is done in a recursive manner in order to provide multiple OSA predictions from each model.
+To that end, we consider a forecasting exercise that explores one-step-ahead (OSA) prediction from $\mathcal{M}_{tv}$ and $\mathcal{M}_{ord}$.
+This is done in a recursive manner to provide multiple OSA predictions from each model.
 We start by defining the initial training data from the beginning of the study until October 2004 (week 4) which contains 200 observations.
 Both models are fitted to the training set and used to predict the next week’s counts (week 1 of November 2004) via the conditional median and mode of the distributions.
-The conditional (or predictive) distributions are Negative-Binomial($\hat{\mu}_{t+1},\hat{\phi}_{t+1}$), or Negative-Binomial($\hat{\mu}_{t+1},\hat{\phi}$) respectively for $\mathcal{M}_{tv}$ and $\mathcal{M}_{ord}$ and $t=200$ at this step.
-Once the prediction is obtained, we add week 1 of November 2004 to the training set, refit both models and gather the new OSA predictions.
+The conditional (or predictive) distributions are $\mbox{NB}(\hat{\mu}_{t+1},\hat{\phi}_{t+1})$ and $\mbox{NB}(\hat{\mu}_{t+1},\hat{\phi})$ respectively for $\mathcal{M}_{tv}$ and $\mathcal{M}_{ord}$ and $t=200$ at this step.
+Once the prediction is obtained, we add week 1 of November 2004 to the training set, refit both models, and gather the new OSA predictions.
 Proceeding until the end of the study period gives a total of 446 predictions from each model.
-Pseudocode describing the steps to this prediction exercise is given in Algorithm 2.
+A pseudocode describing the steps to this prediction exercise is given in Algorithm 2.
 Input: $\boldsymbol{Y}$ observed count trajectory of length $n$
@@
        // train data
-      2. Fit the tv-NBINGARCH model to $\boldsymbol{Y}(s)$;
-      3. From 2, gather the CMLEs $\widehat{\boldsymbol{\theta}}(s)$ and the fitted $\widehat{\boldsymbol{\lambda}}(s)$ and $\widehat{\boldsymbol{\phi}}(s)$ of step $s$;
+      2. Fit the NB tv-DINGARCH model to $\boldsymbol{Y}(s)$;
+      3. From 2, put together the CMLEs $\widehat{\boldsymbol{\theta}}(s)$ and the fitted $\widehat{\boldsymbol{\lambda}}(s)$ and $\widehat{\boldsymbol{\phi}}(s)$ of step $s$;
       4.
 Obtain the OSA mean $\widehat{\mu}_{s+1}=\widehat{\beta}_{0}^{s}+\widehat{\beta}_{1}^{s}Y_{s}+\widehat{\beta}_{2}^{s}\widehat{\mu}_{s}$ and $\widehat{\phi}_{s+1}=\widehat{\alpha}_{0}^{s}+\widehat{\alpha}_{1}^{s}Y_{s}+\widehat{\alpha}_{2}^{s}\widehat{\phi}_{s}$;
-      5A) $\widehat{Y}_{s+1}=\widehat{\mu}_{s+1}$
-       // prediction via mean
-      5B) $\widehat{Y}_{s+1}=\texttt{qbinom}(0.5,\texttt{size}=\widehat{\phi}_{s+1},\texttt{mu}=\widehat{\mu}_{s+1})$
+      5A) Compute the mode of the $\mbox{NB}(\widehat{\mu}_{s+1},\widehat{\phi}_{s+1})$ distribution and denote it by $\widehat{Y}_{s+1}$;
+       // prediction via mode
+      5B) Compute the median of the $\mbox{NB}(\widehat{\mu}_{s+1},\widehat{\phi}_{s+1})$ distribution and denote it by $\widehat{Y}_{s+1}$;
        // prediction via median
@@
 The training data at iteration $s$, $\boldsymbol{Y}(s)$, is incremented and the model is refitted to obtain the OSA forecast $\widehat{Y}_{s+1}$.
-Steps 5A) and 5B) provide alternative prediction methods via the mean or median.
-Algorithm 2 is presented in terms of the $\mathcal{M}_{tv}$ model but works similarly for $\mathcal{M}_{ord}$.
-In this case, $\mathcal{M}_{ord}$ is fitted in step two and the predictions in 5A) and 5B) take in the (fixed) dispersion parameter estimated in step 2.
-By employing the Algorithm with $\mathcal{M}_{tv}$ and $\mathcal{M}_{ord}$, their predictions are summarized via the root mean forecasting error (RMSFE).
-Let $n_{0}$ denote the time point chosen to start the prediction exercise; in our case $n_{0}=200$ the RMSFE of the forecasting step $t$ is
+Steps 5A) and 5B) provide alternative prediction methods via the mode and median.
+Algorithm 2 is presented in terms of the $\mathcal{M}_{tv}$ model but applies similarly for the case of $\mathcal{M}_{ord}$.
+In this case, $\mathcal{M}_{ord}$ is fitted in step two, and the predictions in 5A) and 5B) take in the (fixed) dispersion parameter estimated in step 2.
+By employing the Algorithm with $\mathcal{M}_{tv}$ and $\mathcal{M}_{ord}$, their predictions are summarized via the root mean square forecasting error (RMSFE).
+Let $n_{0}$ denote the time point chosen to start the prediction exercise; in this case $n_{0}=200$.
+The RMSFE of the forecasting step $t$ is
 $\mbox{RMSFE}_{t}=\sqrt{\frac{1}{t-n_{0}}\sum_{s=n_{0}+1}^{t}(Y_{s}-\widehat{Y}_{s})^{2}},$
@@
 Calculation of RMSFE, from the 446 total predictions from $\mathcal{M}_{tv}$ and $\mathcal{M}_{ord}$, yields Figure 4.
-On the left, $\widehat{Y}_{s}$ is the mode of the conditional distribution and on the right it is taken as the median.
+On the left, $\widehat{Y}_{s}$ is the mode of the conditional distribution and on the right, it is taken as the median.
 Prediction by the median yields a smaller prediction error in comparison to the mode, and in both cases the tv-DINGARCH model produces the smallest RMSFE values.
 Figure 4: RMSFE of predictions obtained with the fit of ordinary (dashed lines) and time-varying dispersion (solid lines) INGARCH models to the weekly count of measles in North Rhine-Westphalia, Germany.
-On the left, predicted values are the mode of the predictive distributions, whereas the median is taken as the point on the right.
-This empirical illustration on the weekly number of measles cases in North Rhine-Westphalia, Germany, demonstrated that the time-varying dispersion is a promising and important extension of the ordinary INGARCH processes that achieved improvement in terms of goodness-of-fit and forecasting.
+On the left, predicted values are the mode of the predictive distributions, whereas the median is taken on the right.
+The analysis of the weekly number of measles cases in North Rhine-Westphalia, Germany, demonstrated that the time-varying dispersion is a promising and important extension of the ordinary INGARCH processes that achieved improvement in terms of goodness-of-fit and forecasting.
 6 Conclusions
-We proposed a class of time-varying dispersion INGARCH (tv-DINGARCH) models and explore stochastic properties such as stationarity and ergodicity.
+We proposed a class of time-varying dispersion INGARCH (tv-DINGARCH) models and explored stochastic properties such as stationarity and ergodicity.
 Estimation of parameters was addressed through conditional maximum likelihood estimation (CMLE) and its associated asymptotic theory was established.
@@
 […]
-Let $w$ be a continuous function with compact support on $(0,\infty)^{2}$ and assumed to be bounded $|w|<1$ without loss of generality.
-Consider $k=1$, ${\bf x},{\bf z}\in(0,\infty)^{2}$, and $\epsilon>0$ arbitrary.
-Denote by $f^{mp}_{\boldsymbol{\xi}}(\cdot)$ the probability function of a mixed Poisson distribution with mean $\lambda$ and variance $\lambda+\lambda^{2}/\phi$, where $\boldsymbol{\xi}=(\lambda,\phi)^{\top}$.
+We follow the arguments of Liu (2012) and show explicitly the steps required for the proof based on model (3); further details can be proved along the lines of Liu (2012) and are omitted.
+By working in the same manner as in Liu (2012, pp.106) , we can prove that $\{\boldsymbol{\xi}_{t}\}_{t\in\mathbb{Z}}$ is a weak Feller chain and bounded in probability.
+Therefore, $\{\boldsymbol{\xi}_{t}\}_{t\in\mathbb{Z}}$ has at least one stationary distribution.
+Let $w$ be a continuous function with compact support on $(0,\infty)^{2}$ and assume, without loss of generality, that it is bounded by 1, i.e.
+$|w|<1$.
+Consider $k=1$, ${\bf x}=(x_{1},x_{2}),{\bf z}=(z_{1},z_{2})\in(0,\infty)^{2}$, and $\epsilon>0$ arbitrary.
+Denote by $f^{mp}_{\boldsymbol{\xi};\mathcal{D}}(\cdot)$ the probability function of a mixed Poisson distribution with mean $\lambda$, variance $\lambda+\lambda^{2}/\phi$, and mixing distribution $\mathcal{D}$, where $\boldsymbol{\xi}=(\lambda,\phi)^{\top}$.
 It follows that
 $\left|E(w(\boldsymbol{\xi}_{1})|\boldsymbol{\xi}_{0}={\bf x})-E(w(\boldsymbol{\xi}_{1})|\boldsymbol{\xi}_{0}={\bf z})\right|\leq$
-$\sum_{y=0}^{\infty}|w(\boldsymbol{\tau}+{\bf B}(y,y)^{\top}+{\bf A}\boldsymbol{\bf x})f^{mp}_{\bf x}(y)-w(\boldsymbol{\tau}+{\bf B}(y,y)^{\top}+{\bf A}\boldsymbol{\bf z})f^{mp}_{\bf z}(y)|\leq$
-$\sum_{y=0}^{\infty}f^{mp}_{\bf x}(y)|w(\boldsymbol{\tau}+{\bf B}(y,y)^{\top}+{\bf A}\boldsymbol{\bf x})-w(\boldsymbol{\tau}+{\bf B}(y,y)^{\top}+{\bf A}\boldsymbol{\bf z})|+$ (A-1)
-$\sum_{y=0}^{\infty}|w(\boldsymbol{\tau}+{\bf B}(y,y)^{\top}+{\bf A}\boldsymbol{\bf z})|\,|f^{mp}_{\bf x}(y)-f^{mp}_{\bf z}(y)|.$ (A-2)
+$\sum_{y=0}^{\infty}|w(\boldsymbol{\tau}+{\bf B}(y,y)^{\top}+{\bf A}\boldsymbol{\bf x})f^{mp}_{\bf x;\mathcal{D}}(y)-w(\boldsymbol{\tau}+{\bf B}(y,y)^{\top}+{\bf A}\boldsymbol{\bf z})f^{mp}_{\bf z;\mathcal{D}}(y)|\leq$
+$\sum_{y=0}^{\infty}f^{mp}_{\bf x;\mathcal{D}}(y)|w(\boldsymbol{\tau}+{\bf B}(y,y)^{\top}+{\bf A}\boldsymbol{\bf x})-w(\boldsymbol{\tau}+{\bf B}(y,y)^{\top}+{\bf A}\boldsymbol{\bf z})|+$ (A-1)
+$\sum_{y=0}^{\infty}|w(\boldsymbol{\tau}+{\bf B}(y,y)^{\top}+{\bf A}\boldsymbol{\bf z})|\,|f^{mp}_{\bf x;\mathcal{D}}(y)-f^{mp}_{\bf z;\mathcal{D}}(y)|.$ (A-2)
 We now find an upper bound for the term (A-2).
 Denote by $f^{pois}_{\lambda}(\cdot)$ the probability function of a Poisson distribution with mean $\lambda$.
-From the mixed Poisson stochastic representation, we have that $f^{mp}_{\bf x}(y)=E(f^{pois}_{x_{1}Z_{1}}(y))$, where $Z_{1}$ is the associated latent random variables with distribution depending on $x_{2}$.
-Similar representation holds for $f^{mp}_{\bf z}(y)$ in terms of an associated latent factor $Z_{2}$ with distribution depending on $z_{2}$.
-Using this and the recalling the fact that $|g|<1$, we obtain that
-$\sum_{y=0}^{\infty}|w(\boldsymbol{\tau}+{\bf B}(y,y)^{\top}+{\bf A}\boldsymbol{\bf z})|\,|f^{mp}_{\bf x}(y)-f^{mp}_{\bf z}(y)|\leq\sum_{y=0}^{\infty}|f^{mp}_{\bf x}(y)-f^{mp}_{\bf z}(y)|=$
-$\sum_{y=0}^{\infty}|E(f^{pois}_{x_{1}Z_{1}}(y))-E(f^{pois}_{z_{1}Z_{2}}(y))|\leq E\left(\sum_{y=0}^{\infty}|f^{pois}_{x_{1}Z_{1}}(y)-f^{pois}_{x_{2}Z_{2}}(y)|\right).$ (A-3)
+From the mixed Poisson stochastic representation, we have that $f^{mp}_{\bf x;\mathcal{D}}(y)=E(f^{pois}_{x_{1}Z_{1}}(y))$, where $Z_{1}$ is the associated latent random variables with distribution depending on $x_{2}$.
+Similar representation holds for $f^{mp}_{\bf z;\mathcal{D}}(y)$ in terms of an associated latent factor $Z_{2}$ with distribution depending on $z_{2}$.
+Using this and recalling the fact that $|\omega|<1$, we obtain that
+$\sum_{y=0}^{\infty}|w(\boldsymbol{\tau}+{\bf B}(y,y)^{\top}+{\bf A}\boldsymbol{\bf z})|\,|f^{mp}_{\bf x;\mathcal{D}}(y)-f^{mp}_{\bf z;\mathcal{D}}(y)|\leq\sum_{y=0}^{\infty}|f^{mp}_{\bf x}(y)-f^{mp}_{\bf z}(y)|=$
+$\sum_{y=0}^{\infty}|E(f^{pois}_{x_{1}Z_{1}}(y))-E(f^{pois}_{z_{1}Z_{2}}(y))|\leq E\left(\sum_{y=0}^{\infty}|f^{pois}_{x_{1}Z_{1}}(y)-f^{pois}_{z_{1}Z_{2}}(y)|\right).$ (A-3)
 By using inequality (A.1) from Wang et al.
-(2014) (see also Liu (2012)), we have that $\sum_{y=0}^{\infty}|f^{pois}_{x_{1}Z_{1}}(y)-f^{pois}_{x_{2}Z_{2}}(y)|\leq 2\left(1-\exp\{-|x_{1}Z_{1}-x_{2}Z_{2}|\}\right)$.
+(2014) (see also Liu (2012)), we have that $\sum_{y=0}^{\infty}|f^{pois}_{x_{1}Z_{1}}(y)-f^{pois}_{z_{1}Z_{2}}(y)|\leq 2\left(1-\exp\{-|x_{1}Z_{1}-z_{1}Z_{2}|\}\right)$.
 Hence, (A-1) is bounded above by
-$E\left(\sum_{y=0}^{\infty}|f^{pois}_{x_{1}Z_{1}}(y)-f^{pois}_{x_{2}Z_{2}}(y)|\right)\leq 2E\left(1-e^{-|x_{1}Z_{1}-x_{2}Z_{2}|}\right)\leq$
-$2E\left(1-e^{-|x_{1}-x_{2}|(Z_{1}+Z_{2})}\right)\leq 2E\left(|x_{1}-x_{2}|(Z_{1}+Z_{2})\right)=2|x_{1}-x_{2}|\leq 2\|{\bf x}-{\bf z}\|_{p},$
+$E\left(\sum_{y=0}^{\infty}|f^{pois}_{x_{1}Z_{1}}(y)-f^{pois}_{z_{1}Z_{2}}(y)|\right)\leq 2E\left(1-e^{-|x_{1}Z_{1}-z_{1}Z_{2}|}\right)\leq$
+$2E\left(1-e^{-|x_{1}-z_{1}|(Z_{1}+Z_{2})}\right)\leq 2E\left(|x_{1}-z_{1}|(Z_{1}+Z_{2})\right)=2|x_{1}-z_{1}|\leq 2\|{\bf x}-{\bf z}\|_{p},$
 for $p\in[1,\infty]$, where we have used in the second inequality the fact that $|ab-cd|\leq|a-c|(b+d)$ for $a,b,c,d>0$.
@@
 108).
-The upper bound for the term given in (A-1) follows exactly as discussed in Liu (2012): choose $\epsilon^{\prime}>0$ and $\eta>0$ small enough such as $\epsilon^{\prime}+\dfrac{8\eta}{1-\|{\bf A}\|_{p}}<\epsilon$ and $\|{\bf x}-{\bf z}\|_{p}<\eta$ implying $|w({\bf x})-w({\bf z})|<\epsilon^{\prime}$, with $p\in[1,\infty]$.
-In this part, we are using the fact that $\|{\bf A}\|_{p}<1$ for some $p\in[1,\infty]$, which follows from the assumption that there exists $p$ such that $\|{\bf A}\|_{p}+2^{1-1/p}\|{\bf B}\|_{p}<1$.
+The upper bound for the term given in (A-1) follows exactly as discussed in Liu (2012): choose $\epsilon^{\prime}>0$ and $\eta>0$ small enough such as $\epsilon^{\prime}+\dfrac{8\eta}{1-\|{\bf A}\|_{p}}<\epsilon$ and $\|{\bf x}-{\bf z}\|_{p}<\eta$ implies $|w({\bf x})-w({\bf z})|<\epsilon^{\prime}$, with $p\in[1,\infty]$.
+In this part, we are using that $\|{\bf A}\|_{p}<1$ for some $p\in[1,\infty]$.
 By combining the above results, we obtain that $\left|E(w(\boldsymbol{\xi}_{1})|\boldsymbol{\xi}_{0}={\bf x})-E(w(\boldsymbol{\xi}_{1})|\boldsymbol{\xi}_{0}={\bf z})\right|<\epsilon^{\prime}+2\|{\bf x}-{\bf z}\|_{p}$.
-For general $k\geq 2$, the result follows by using mathematical induction, exactly as done in Chapter 4 from Liu (2012), and therefore it is omitted.
-With the e-chain property established for the bivariate process $\{\boldsymbol{\xi}_{t}\}$ and under the assumption of existence of $p\in[1,\infty]$ such that $\|{\bf A}\|_{p}+2^{1-1/p}\|{\bf B}\|_{p}<1$, following the same steps that proofs of Proposition 4.2.1(b) and Proposition 4.3.1 from Liu (2012), we obtain the desired result.
+For general $k\geq 2$, the result follows by using mathematical induction, exactly as done in Chapter 4 of Liu (2012), and therefore it is omitted.
+With the e-chain property established for the bivariate process $\{\boldsymbol{\xi}_{t}\}_{t\in\mathbb{Z}}$, we show part (b) by following similar arguments of the proofs of Proposition 4.2.1(b) and Proposition 4.3.1 of Liu (2012) to establish the geometric moment contraction (GMC) of $\{\boldsymbol{\xi}_{t}\}_{t\in\mathbb{Z}}$.
+Denote by $(Y_{t},\boldsymbol{\xi}_{t})$ and $(Y_{t}^{(0)},\boldsymbol{\xi}_{t}^{(0)})$ the process under the initial value $\boldsymbol{\xi}_{0}=\boldsymbol{\xi}=(\lambda,\phi)^{\top}$ and $\boldsymbol{\xi}_{0}=\boldsymbol{\xi}^{(0)}=(\lambda^{(0)},\phi^{(0)})^{\top}$, respectively.
+It follows that
+$E\left(\|\boldsymbol{\xi}_{1}-\boldsymbol{\xi}_{1}^{(0)}\|_{p}\right)$ $=$ $E\left(\|(\boldsymbol{\tau}+{\bf A}\boldsymbol{\xi}+{\bf B}(Y_{0},Y_{0})^{\top})-(\boldsymbol{\tau}+{\bf A}\boldsymbol{\xi}^{(0)}+{\bf B}(Y_{0}^{(0)},Y_{0}^{(0)})^{\top})\|_{p}\right)$
+$\leq$ $\|{\bf A}\|_{p}\|\boldsymbol{\xi}-\boldsymbol{\xi}^{(0)}\|_{p}+E\left(\|{\bf B}(Y_{0}-Y_{0}^{(0)},Y_{0}-Y_{0}^{(0)})^{\top}\|_{p}\right)$
+$=$ $\|{\bf A}\|_{p}\|\boldsymbol{\xi}-\boldsymbol{\xi}^{(0)}\|_{p}+\|{\bf B}\|_{p}E\left(|Y_{0}-Y_{0}^{(0)}|\right).$
+Further, we have that $E(|Y_{0}-Y_{0}^{(0)}|)=|\lambda-\lambda^{(0)}|$; for instance, see proof of Thm 1 by Christou and Fokianos (2014). Hence,
+$E\left(\|\boldsymbol{\xi}_{1}-\boldsymbol{\xi}_{1}^{(0)}\|_{p}\right)$$\leq$ $\|{\bf A}\|_{p}\|\boldsymbol{\xi}-\boldsymbol{\xi}^{(0)}\|_{p}+\|{\bf B}\|_{p}|\lambda-\lambda^{(0)}|$
+$\leq$ $\|{\bf A}\|_{p}\|\boldsymbol{\xi}-\boldsymbol{\xi}^{(0)}\|_{p}+\|{\bf B}\|_{p}\|\boldsymbol{\xi}-\boldsymbol{\xi}^{(0)}\|_{1}$
+$\leq$ $\|{\bf A}\|_{p}\|\boldsymbol{\xi}-\boldsymbol{\xi}^{(0)}\|_{p}+2^{1-1/p}\|{\bf B}\|_{p}\|\boldsymbol{\xi}-\boldsymbol{\xi}^{(0)}\|_{p}$
+$\leq$ $(\|{\bf A}\|_{p}+2^{1-1/p}\|{\bf B}\|_{p})\|\boldsymbol{\xi}-\boldsymbol{\xi}^{(0)}\|_{p},$
+where we have used that $\|\boldsymbol{\xi}-\boldsymbol{\xi}^{(0)}\|_{1}\leq 2^{1-1/p}\|\boldsymbol{\xi}-\boldsymbol{\xi}^{(0)}\|_{p}$.
+Since $\|{\bf A}\|_{p}+2^{1-1/p}\|{\bf B}\|_{p}<1$, the GMC is satisfied, which concludes the proof.
+$\square$
 A-2 Proof of Theorem 2.2
-Let $\{Y_{t}\}_{t\in\mathbb{N}}$ be a first-order negative binomial INGARCH process.
-Under condition $\|{\bf A}\|_{1}+\|{\bf B}\|_{1}<1$, Theorem 2.1 holds (for $p=1$) and gives us that $\{Y_{t}\}_{t\in\mathbb{N}}$ has a unique stationary and ergodic solution.
+Let $\{Y_{t}\}_{t\in\mathbb{N}}$ be a first-order negative binomial tv-DINGARCH process.
+Under condition $\|{\bf A}\|_{1}+\|{\bf B}\|_{1}<1$, Theorem 2.1 holds (for $p=1$) and gives us that $\{\boldsymbol{\xi}_{t}\}_{t\in\mathbb{Z}}$ has a stationary and ergodic solution.
 Moreover, this process is a Markov chain.
-Therefore, we will now show that Tweedie’s criterion is satisfied for a test function to show the finiteness of arbitrary moments stated in the lemma (Meyn and Tweedie, 1993).
+Therefore, we need to show that Tweedie’s criterion (for instance, see Chapter 15 from Meyn and Tweedie (1993)) is satisfied for an appropriate test function to prove the finiteness of arbitrary moments.
 Consider the test function $V(x,z)=1+x^{k}+z^{k}$, for $x,z>0$ and arbitrary $k\in\mathbb{N}$. It follows that
@@
 $E(Y^{d})=\lambda\left\{1+\sum_{j=1}^{d-1}\left[\binom{d-1}{j}+\dfrac{1}{\phi}\binom{d-1}{j-1}\right]E(Y^{j})\right\},$ (A-5)
-with $E(Y)=\lambda$.
-From (A-5), we immediately obtain that $E(Y^{d})=\lambda^{d}(1+1/\phi)+O(\lambda^{d-1})$ as $\lambda\rightarrow\infty$. Using this result, it follows that
+with $E(Y)=\lambda$, which yields that $E(Y^{d})=\lambda^{d}(1+1/\phi)+O(\lambda^{d-1})$ as $\lambda\rightarrow\infty$. Using this result, it follows that
 $E\left((\beta_{0}+\beta_{1}Y+\beta_{2}\lambda)^{k}\right)$ $=$ $\sum_{j=0}^{k}\binom{k}{j}(\beta_{0}+\beta_{2}\lambda)^{k-j}\beta_{1}^{j}E(Y^{j})=\sum_{j=0}^{k}\binom{k}{j}(\beta_{2}\lambda)^{k-j}\beta_{1}^{j}\lambda^{j}+O(\lambda^{k-1})$
@@
 $E\left((\alpha_{0}+\alpha_{1}Y+\alpha_{2}\phi)^{k}\right)=(\alpha_{1}\lambda+\alpha_{2}\phi)^{k}+O(\lambda^{k-1}).$
-Hence, we can express the conditional expectation (A-4) as
+Hence, (A-4) becomes
 $E\left(V(\lambda_{t},\phi_{t})|\lambda_{t-1}=\lambda,\phi_{t-1}=\phi\right)=V(\lambda,\phi)\dfrac{1+\lambda^{k}(\beta_{1}+\beta_{2})^{k}+(\alpha_{1}\lambda+\alpha_{2}\phi)^{k}+O(\lambda^{k-1})}{1+\lambda^{k}+\phi^{k}}.$ (A-6)
@@
 as $r\rightarrow\infty$, where the last inequality follows from lemma’s assumption implying that $\beta_{1}+\beta_{2}<1$ and $\alpha_{1}+\alpha_{2}<1$.
-Further, the remaining term converges to 0 as $r\rightarrow\infty$: $\dfrac{O(\lambda^{k-1})}{1+\lambda^{k}+\phi^{k}}=O(r^{k-1})/O(r^{k})=O(r^{-1})$.
-Therefore, there exist real constants $\kappa_{1}\in(0,1)$, $\kappa_{2}>0$, and $L>0$ such that (A-4) is bounded above as follows:
+Furthermore, the remaining term converges to 0 as $r\rightarrow\infty$ because $\dfrac{O(\lambda^{k-1})}{1+\lambda^{k}+\phi^{k}}=O(r^{k-1})/O(r^{k})=O(r^{-1})$.
+Therefore, there exist real constants $\kappa_{1}\in(0,1)$, $\kappa_{2}>0$, and $L>0$ such that (A-4) is bounded from above as follows:
 $E\left(V(\lambda_{t},\phi_{t})|\lambda_{t-1}=\lambda,\phi_{t-1}=\phi\right)\leq(1-\kappa_{1})V(\lambda,\phi)+\kappa_{2}I\{(\lambda,\phi)\in G\},$
@@
 Since the $l$-th moment of $\lambda_{t}$ is finite for arbitrary $l\in\mathbb{N}$, we obtain that the unconditional $k$-th moment of $Y_{t}$ is finite for all $k\geq 1$ as well.
+$\square$
 Lemma A-1.
+Under the conditions of Theorem 3.1
+$\sup\limits_{\boldsymbol{\theta}\in\mathcal{B}}\Big{|}\frac{1}{n}\ell(\boldsymbol{\theta})-\frac{1}{n}\tilde{\ell}(\boldsymbol{\theta})\Big{|}\overset{a.s.}{\longrightarrow}0\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \mbox{as}\leavevmode\nobreak\ n\rightarrow\infty$
+following the notation from Subsection 3.1.
+Proof.
+Let $\boldsymbol{\xi}_{t}=(\lambda_{t},\phi_{t})^{\top}$ denote the stationary and ergodic solution of model (3).
+In addition, let $\tilde{\boldsymbol{\xi}}_{t}=(\tilde{\lambda}_{t},\tilde{\phi}_{t})^{\top}$ the process obtained with some starting value $\tilde{\boldsymbol{\xi}}_{0}=(\tilde{\lambda}_{0},\tilde{\phi}_{0})^{\top}$.
+Then, arguing as in Christou and Fokianos (2014, Lemma A.2) and using that $\|{\bf A}\|_{1}<1$.
+we obtain that
+$\sup\limits_{\boldsymbol{\theta}\in\mathcal{B}}\|\boldsymbol{\xi}_{t}-\tilde{\boldsymbol{\xi}}_{t}\|_{1}\leq C\|{\bf A}\|^{t}_{1},\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \forall\leavevmode\nobreak\ t,$
+almost surely from the compactness of $\mathcal{B}$, for some constant $C>0$.
+Now, consider the differences $\ell_{t}(\boldsymbol{\theta})-\tilde{\ell}_{t}(\boldsymbol{\theta})$ following the notation introduced by (5) and (6).
+By grouping terms we obtain the following inequalities using repeatedly the facts that for any $x,y>0$ it holds that $|\log(x/y)|\leq|x-y|/\min(x,y)$, $\lambda_{t}>\beta_{0}$, $\phi_{t}>\alpha_{0}$ and $\log x<x-1$, for all $x>0$.
+$\bigg{|}y_{t}\Bigl{[}\log\lambda_{t}-\log(\lambda_{t}+\phi_{t})\Bigr{]}-y_{t}\Bigr{[}\log\tilde{\lambda}_{t}-\log(\tilde{\lambda}_{t}+\tilde{\phi}_{t})\Bigl{]}\bigg{|}$ $\leq 2y_{t}\frac{\|\xi_{t}-\tilde{\xi}_{t}\|_{1}}{\beta_{0}},$
+$\bigg{|}\phi_{t}\Bigl{[}\log\phi_{t}-\log(\lambda_{t}+\phi_{t})\Bigr{]}-\tilde{\phi}_{t}\Bigr{[}\log\tilde{\phi}_{t}-\log(\tilde{\lambda}_{t}+\tilde{\phi}_{t})\Bigl{]}\bigg{|}$ $\leq\|\xi_{t}-\tilde{\xi}_{t}\|_{1}\Bigl{\{}\Bigl{[}(\frac{2}{\alpha_{0}}+1)+\phi_{t}+\lambda_{t}+2\|\xi_{t}-\tilde{\xi}_{t}\|_{1}\Bigr{]}\Bigr{\}},$
+$\Bigr{|}\log\Gamma(y_{t}+\phi_{t})-\log\Gamma(y_{t}+\tilde{\phi}_{t})\Bigl{|}\leq(\frac{2}{\alpha_{0}}+1)(y_{t}+\phi_{t}+1)\|\xi_{t}-\tilde{\xi}_{t}\|_{1},$
+$\Bigr{|}\log\Gamma(\phi_{t})-\log\Gamma(\tilde{\phi}_{t})\Bigl{|}\leq(\frac{2}{\alpha_{0}}+1)(\phi_{t}+1)\|\xi_{t}-\tilde{\xi}_{t}\|_{1}.$
+For the last two inequalities, we employ Stirling’s approximation to the Gamma function. In conclusion
+$\sup\limits_{\boldsymbol{\theta}\in\mathcal{B}}\Big{|}\frac{1}{n}\ell(\boldsymbol{\theta})-\frac{1}{n}\tilde{\ell}(\boldsymbol{\theta})\Big{|}$$\leq$ $\frac{C}{n}\sum_{t=1}^{n}\|\boldsymbol{A}\|^{t}_{1}W_{t},$
+where $W_{t}$ is a positive linear combination of $(Y_{t},\lambda_{t},\phi_{t})$.
+Theorem 2.2 implies that $E|W_{t}|^{k}<\infty$, for all $k\in\mathbb{N}$.
+But using Markovs inequality for every $\epsilon>0$ we have that
+$\sum_{t=1}^{\infty}P\Big{(}\|\boldsymbol{A}\|^{t}_{1}W_{t}>\epsilon\Big{)}\leq\displaystyle\sum_{t=1}^{\infty}\frac{\|\boldsymbol{A}\|^{kt}_{1}E(W_{t}^{k})}{\epsilon^{k}}<\infty.$
+Hence, $\|\boldsymbol{A}\|^{t}_{1}W_{t}\overset{a.s.}{\longrightarrow}0$, as $t\rightarrow\infty$ and therefore the proof is concluded. ∎
+A-3 Proof of Lemma 3.1
+We have that $|S_{1t}|\leq|y_{t}/\lambda_{t}-1|\leq y_{t}/\beta_{0}+1$.
+Furthermore, $\partial\lambda_{t}/\partial\beta_{i}$ is a linear combination of $y_{t-1},\ldots,y_{1}$ and $\lambda_{1}$, for $i=0,1,2$.
+Using these results and under condition $\|{\bf A}\|_{1}+\|{\bf B}\|_{1}<1$, Theorem 2.2 gives that the score function associated with $\beta_{0},\beta_{1}$ and $\beta_{2}$ is integrable.
+We now study the score function associated with $\alpha_{0}$, $\alpha_{1}$ and $\alpha_{2}$.
+Note that $\left|-\dfrac{y_{t}-\lambda_{t}}{\lambda_{t}+\phi_{t}}\right|\leq\dfrac{y_{t}}{\lambda_{t}+\phi_{t}}+1\leq\dfrac{y_{t}}{2}\left(\dfrac{1}{\beta_{0}}+\dfrac{1}{\alpha_{0}}\right)+1$, where we have used that fact that $\dfrac{1}{a+b}\leq\dfrac{1}{2}\left(\dfrac{1}{a}+\dfrac{1}{b}\right)$ for $a,b>0$.
+Also, by using that $\dfrac{1}{2x}<\log x-\Psi(x)<\dfrac{1}{x}$ for $x>0$ (Alzer, 1997), we obtain that $0<\log\phi_{t}-\Psi(\phi_{t})<\phi_{t}^{-1}<\alpha_{0}^{-1}$.
+Similarly, it follows that $|\Psi(\lambda_{t}+\phi_{t})-\log(\lambda_{t}+\phi_{t})|<(\lambda_{t}+\phi_{t})^{-1}\leq\dfrac{1}{2}\left(\dfrac{1}{\beta_{0}}+\dfrac{1}{\alpha_{0}}\right)$.
+Therefore, $|S_{2t}|$ is bounded above by a linear combination of $y_{t}$.
+Moreover, $\dfrac{\partial\phi_{t}}{\partial\alpha_{j}}$ is a linear combination of $y_{t-1},\ldots,y_{1}$ and $\phi_{1}$, for $j=0,1,2$.
+Again, an application of Theorem 2.2 gives that the score function associated with $\alpha_{0}$, $\alpha_{1}$ and $\alpha_{2}$ is also integrable.
+Lemma A-2.
 Let $\mathcal{B}$ as defined in (11).
@@
 Proof.
-We will show that the result holds for the third derivative of $\ell(\cdot)$ with respect to $\beta_{1}$ and also $\alpha_{1}$.
-The remaining cases follow the very same steps as these two and therefore are omitted.
+We will show that the result holds for the third derivative of $\ell(\cdot)$ with respect to $\beta_{1}$ and also $\alpha_{1}$. The remaining cases are similar and therefore are omitted.
 The third derivative of $\ell_{t}(\boldsymbol{\theta})$ with respect to $\beta_{1}$ is given by $\dfrac{\partial^{3}\ell_{t}(\boldsymbol{\theta})}{\partial\beta_{1}^{3}}=\dfrac{\partial^{2}S_{1t}}{\partial\phi_{t}^{2}}\left(\dfrac{\partial\lambda_{t}}{\partial\beta_{1}}\right)^{3}$, where
@@
 (2009) (proof of Lemma 3.4 of their Supplementary Material).
-The case for $\alpha_{1}$ is more involving.
+The case for $\alpha_{1}$ is more involved.
 We have that $\dfrac{\partial^{3}\ell_{t}(\boldsymbol{\theta})}{\partial\alpha_{1}^{3}}=\dfrac{\partial^{2}S_{2,t}}{\partial\phi_{t}^{2}}\left(\dfrac{\partial\phi_{t}}{\partial\alpha_{1}}\right)^{3}$, where
@@
 Inequality (15) from Guo and Qi (2013) gives that $|\Psi^{\prime\prime}(x)|<\dfrac{1}{x}+\dfrac{2}{x^{3}}$, for $x>0$.
-Hence, it follows that $|\Psi^{\prime\prime}(y_{t}+\phi_{t})|<\dfrac{1}{(y_{t}+\phi_{t})^{2}}+\dfrac{2}{(y_{t}+\phi_{t})^{3}}\leq\dfrac{1}{\phi_{t}^{2}}+\dfrac{2}{\phi_{t}^{3}}\leq\dfrac{1}{\alpha_{0}^{2}}+\dfrac{2}{\alpha_{0}^{3}}\leq\dfrac{1}{\alpha_{0,inf}^{2}}+\dfrac{2}{\alpha_{0,inf}^{3}}$.
-Similarly, $|\Psi^{\prime\prime}(\phi_{t})|<\dfrac{1}{\alpha_{0,inf}^{2}}+\dfrac{2}{\alpha_{0,inf}^{3}}$.
+Hence, it follows that $|\Psi^{\prime\prime}(y_{t}+\phi_{t})|<\dfrac{1}{(y_{t}+\phi_{t})^{2}}+\dfrac{2}{(y_{t}+\phi_{t})^{3}}\leq\dfrac{1}{\phi_{t}^{2}}+\dfrac{2}{\phi_{t}^{3}}\leq\dfrac{1}{\alpha_{0}^{2}}+\dfrac{2}{\alpha_{0}^{3}}\leq\dfrac{1}{\alpha_{0,low}^{2}}+\dfrac{2}{\alpha_{0,low}^{3}}$.
+Similarly, $|\Psi^{\prime\prime}(\phi_{t})|<\dfrac{1}{\alpha_{0,low}^{2}}+\dfrac{2}{\alpha_{0,low}^{3}}$.
 The modulus of the remaining terms in (A-7) are bounded above (as done for the third derivative involving $\beta_{1}$) by a linear function of $y_{t}$ with positive coefficients do not depend on $\boldsymbol{\theta}$.
@@
 ∎
-A-3 Proof of Theorem 3.1
-We will show that the Conditions (A1), (A2), and (A3) from Lemma 3.1 by Jensen and Rahbek (2004) are satisfied, which give us the desired results stated in the theorem.
+A-4 Proof of Theorem 3.1
+We will show that the Conditions (A1), (A2), and (A3) from Lemma 3.1 by Jensen and Rahbek (2004) (see also Fokianos et al.
+(2009) and Christou and Fokianos (2014)) are satisfied, which yield the theorem.
 As argued before for the general case, we use Proposition 3.1 and the Central Limit Theorem for martingale difference to establish the weak convergence involving the score function in (8).
-The existence and finiteness of the matrix $\boldsymbol{\Omega}_{1}(\boldsymbol{\theta})$ follow from Theorem 2.2 that provides the finiteness of the moments of all orders for the linear NB tv-DINARCH process under the assumption that $\|{\bf A}\|_{1}+\|{\bf B}\|_{1}<1$, which is in force.
+The existence and finiteness of the matrix $\boldsymbol{\Omega}_{1}(\boldsymbol{\theta})$ follow from Theorem 2.2 that provides the finiteness of the moments of all orders for the linear NB tv-DINARCH process under the assumption that $\|{\bf A}\|_{1}+\|{\bf B}\|_{1}<1$.
 Therefore, Condition (A1) holds.
-Moreover, under the assumption $\|{\bf A}\|_{1}+\|{\bf B}\|_{1}<1$, we obtain from Theorem 2.1 that $\{Y_{t}\}_{t\in\mathbb{N}}$ is a stationary and ergodic sequence.
-Then, the Law of Large Numbers for stationary and ergodic sequences gives us that (6) is valid, with the existence and finiteness of the matrix $\boldsymbol{\Omega}_{2}(\boldsymbol{\theta})$ being ensured again by Theorem 2.2.
-That is, Condition (A2) from Lemma 3.1 by Jensen and Rahbek (2004) is satisfied.
-Finally, Condition (A3) has been established in Lemma A-1.
+Moreover, we obtain from Theorem 2.1 that $\{Y_{t}\}_{t\in\mathbb{Z}}$ is a stationary and ergodic sequence.
+Then, the Law of Large Numbers for stationary and ergodic sequences gives us that (9) holds true, with the existence and finiteness of the matrix $\boldsymbol{\Omega}_{2}(\boldsymbol{\theta})$ being ensured again by Theorem 2.2.
+We prove next that $\boldsymbol{\Omega}_{1}(\boldsymbol{\theta})$ is non-singular.
+Suppose that it is singular.
+Then there exist vectors $\nu_{1}$, $\nu_{2}$ both in $\mathbb{R}^{3}$ such that the six-dimensional vector $\nu=(\nu_{1}^{\top},\nu_{2}^{\top})^{\top}$ satisfies $\nu_{1}^{\top}{\partial\lambda_{t}}/{\partial\boldsymbol{\theta}_{1}}+\nu_{2}^{\top}{\partial\phi_{t}}/{\partial\boldsymbol{\theta}_{2}}=0$, a.s for $t\geq 1$.
+But both ${\partial\lambda_{t}}/{\partial\boldsymbol{\theta}_{1}}$ and ${\partial\phi_{t}}/{\partial\boldsymbol{\theta}_{2}}$ are stationary so this holds for all $t$.
+However,
+$\frac{\partial\lambda_{t}}{\partial\boldsymbol{\theta}_{1}}$ $=(1,Y_{t-1},\lambda_{t-1})^{\top}+\beta_{2}\frac{\partial\lambda_{t-1}}{\partial\boldsymbol{\theta}_{1}},$
+$\frac{\partial\phi_{t}}{\partial\boldsymbol{\theta}_{2}}$ $=(1,Y_{t-1},\phi_{t-1})^{\top}+\alpha_{2}\frac{\partial\phi_{t-1}}{\partial\boldsymbol{\theta}_{2}}.$
+The matrix $\boldsymbol{\Omega}_{1}(\boldsymbol{\theta})$ is block-diagonal.
+Therefore it is singular if and only if $\nu_{1}^{\top}(1,Y_{t-1},\lambda_{t-1})^{\top}+\nu_{2}^{\top}(1,Y_{t-1},\phi_{t-1})^{\top}=0$ a.s for all $t\geq 1$.
+But this cannot hold because of the constant term 1 and the fact that $Y_{t}$, $\lambda_{t}$ and $\phi_{t}$ are non-zero, for some $t$.
+That is, Condition (A2) from Lemma 3.1 by Jensen and Rahbek (2004) is satisfied. Finally, Condition (A3) has been established in Lemma A-2.
 Acknowledgments
-H. Ombao would like to acknowledge support by KAUST Research Fund. L.S.C. Piancastelli wishes to acknowledge the financial support of Science Foundation Ireland.
+We would like to thank the Editor and two reviewers for their careful reading and constructive comments that improved the presentation.
+H.
+Ombao would like to acknowledge support from the KAUST Research Fund.
+L.S.C.
+Piancastelli wishes to acknowledge the financial support from the Science Foundation Ireland.
+Declarations
+Conflict of interest: The authors have no conflicts of interest to declare.